### v1

# Extracting the singularity of the logarithmic partition function

Abstract
In principle, the logarithm of the partition function and its derivatives can be used to rigorously determine the nature of phase transitions. However, in practice, this approach is often ineffective in numerical calculations, since resolving the singularities of the logarithmic partition function requires extremely high-precision data—especially when distinguishing between weakly first-order phase transitions and deconfined quantum critical points. Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the leading volume-law term. This is achieved by computing the Rényi thermal entropy (RTE) and its derivative (DRTE). We have derived the scaling relations for RTE and DRTE, which successfully determine the phase transition points and critical exponents through data collapse in various quantum many-body systems. Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in $J\text{-}Q$ models, which were initially thought to be continuous deconfined quantum phase transitions. The discontinuity of these transitions is challenging to detect using traditional local physical quantities. This work provides a new and powerful paradigm for studying quantum many-body phase transitions.
Introduction.— Quantum phase transitions, spontaneous symmetry breaking, and the universality classes of critical phenomena are central concepts in condensed matter and statistical physics [1, 2, 3]. The type of phase transition is well defined by the singularity of free energy and its derivative. For a first-order phase transition, the first-order derivative of free energy is discontinuous. Otherwise, if its higher-order derivative is discontinuous, it is named a continuous phase transition. In principle, if the partition function can be calculated exactly, the phase transition will be fixed strictly according to the singularity. However, in practical numerical simulations, directly detecting such singularities through the free energy and its derivatives is difficult, as it requires extremely high-precision data. This challenge is particularly pronounced when distinguishing between weakly first-order phase transitions and deconfined quantum critical points  [4, 5, 6, 7, 8, 9, 10]. A major reason for this difficulty is that the singular contribution is typically overshadowed by a much larger, non-singular term that scales with the system volume. How to extract the singularity remains a significant challenge in theoretically computational physics.
Considering a quantum phase transition ($\beta\rightarrow\infty$), the singularity of free energy can be replaced by focusing on the logarithmic partition function. According to the finite-size scaling hypothesis [11, 12], an observable which is singular at $g_{c}$ (the parameter $g$ of Hamiltonian at the phase transition point) scales with the system size close to $g_{c}$ as a power of $L$ multiplied by a non-singular function of the ratio $\xi/L$, where $\xi$ is the correlation length and $L$ is the system size. Any singular quantity (not necessarily divergent) should thus be of the form:
$Q(\Delta g,L)=L^{\sigma}f(\Delta gL^{1/\nu})$ (1)
where $\Delta g=g-g_{c}$, $\nu$ is the critical exponent of the correlation length and $\sigma$ is an undetermined power. Since the singular part of the free energy is scale-invariant ($\sigma=0$[13, 11], the logarithmic partition function can be written in the following form:
$\ln Z(g)=a(g)L^{d}+f(\Delta gL^{1/\nu})$ (2)
where $Z$ is a partition function, $d$ is the spatial dimension and $a(g)L^{d}$ represents the leading term obeying a volume law. Here, we see that this dominant volume law term often masks the singular contribution of $f$, which explains the difficulty in numerically detecting the phase transition through such singularities.
An approach to extract the singular term $f$ is to construct a quantity in which the volume-law contributions cancel out, leaving only the singular part. Inspired by the method used to calculate topological entanglement entropy [14, 15], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\rightarrow\infty$) isolates the singular term $f$, with the dominant volume-law contributions canceled out. Specifically, we argue that Eq. (2) can be rewritten as $\ln Z(2\beta)=2a(g)L^{d}+f$ and $2\ln Z(\beta)=2a(g)L^{d}+2f$. Coincidentally, we find that this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [16, 17, 18, 19, 20, 21, 22]:
$S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$ (3)
This definition can be generalized to an arbitrary order $n$, leading to the $n$-th order RTE: $S^{(n)}=\frac{1}{1-n}\ln[Z(n\beta)/Z(\beta)^{n}]$. For simplicity, we focus on the 2nd-order RTE ( i.e., $n=2$) throughout this paper.
In this work, assisted by our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [23, 24, 25, 26, 27, 28, 29], we demonstrate the validity of our conjecture above: the RTE is indeed a dimensionless quantity that captures only the singular contribution. Furthermore, we observe a more pronounced singularity in the derivative of the RTE (DRTE) near the quantum critical point (QCP). We propose universal scaling relations for both the RTE and DRTE, which are confirmed by data collapse analyses performed near (2+1)D QCPs belonging to the O($N$) universality classes, with $1\leq N\leq 3$. More interestingly, since the leading non-singular term of $\ln Z$ is canceled out in the DRTE, it becomes a particularly powerful tool for detecting weakly first-order phase transitions even in normal sizes, which are notoriously difficult to resolve using conventional numerical approaches.
Models and method.— We investigate quantum spin models with continuous QCPs in the (2+1)D O($N$) universality classes, where $1\leq N\leq 3$. Additionally, we explore first-order transitions, with a particular focus on the challenging weakly first-order phase transitions (e.g., some controversial deconfined quantum critical points). Detailed descriptions of the model Hamiltonians will be provided in the following sections.
As shown by the definitions above, the 2nd-order Rényi thermal entropy can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$. The partition functions can be calculated using the bipartite reweight-annealing algorithm [23, 24] in conjunction with stochastic series expansion (SSE) QMC simulations [30, 12, 31, 32, 33, 34, 35]. Furthermore, following a procedure similar to that used in calculating the derivative of Rényi entanglement entropy and negativity [23, 36], the derivative of the RTE with respect to a general parameter $J$ can be written as:
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
In this expression, the first average is evaluated on the ensemble corresponding to $Z(n\beta)$, while the second is based on $Z(\beta)$. In our simulations, the inverse temperature $\beta$ is set to be linearly proportional to the lattice size $L$ (i.e., $\beta=L$) in order to approach the quantum critical regime. Further methodological details are provided in the Supplementary Materials (SM).
Finite-size scaling form.— Considering a zero-temperature system and assuming the ground-state degeneracy is $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$+$f(\Delta gL^{1/\nu})$, where $f$ represents the critical corrections. As shown in Fig.1, $S^{(2)}$ can be used to probe both the degeneracy of discrete symmetry-breaking and the tower of excited states (TOS) associated with continuous symmetry-breaking.
More interestingly, as shown in Fig. 1, we observe that the curves for different system sizes cross at the critical point, which implies that $S^{(2)}$ is indeed a dimensionless quantity, consistent with our argument in the introduction. Near the critical point, the logarithmic partition function can be decomposed into a non-singular part $a(J)L^{d}$ and a dimensionless singular term $f(\Delta gL^{1/\nu})$. The non-singular contribution is the leading term, which makes it difficult to extract critical information directly from the partition function. As shown in Fig. 1, the crossings of curves for different sizes indicate that the RTE is independent of the system size. This provides strong numerical support for our argument that the leading volume-law contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel out in the definition of RTE, leaving only the dimensionless singular part. Additionally, we observe that the DRTE exhibits a stronger divergence near the QCP.
According to the finite-size scaling hypothesis  [11, 12], any singular quantity near a critical point should obey a scaling form as given in Eq.(1). In particular, the dimensionless RTE ($\sigma=0$), is expected to follow the scaling form:
$S^{(n)}(\Delta g,L)=f(\Delta gL^{1/\nu}).$ (5)
From Eq. (5), we can readily derive the scaling behavior of the DRTE:
$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=f^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
where $f^{\prime}(x)$ is $\partial f(x)/\partial x$ and $x$ is a general parameter.
Figure 1: (a) The 2nd-order RTE and (c) its derivative, and (b) and (d) the data collapse analysis of the RTE and DRTE near the QCP of the bilayer Ising-Heisenberg model (with (2+1)D Ising universality class). We set $J=3.045$ and take $J_{z}$ as the tuning parameter. (e–h) Data of the dimerized anisotropic Heisenberg model with $\Delta=0.9$ (with (2+1)D O(2) universality class). We set $J_{2}=2.1035$, and take $J_{1}$ as the tuning parameter. (i–m) Data of the dimerized isotropic Heisenberg model with $\Delta=1$ (with (2+1)D O(3) universality class). We set $J_{2}=1$, and take $J_{1}$ as the tuning parameter. To achieve good data collapse, a subleading correction is also included: $Q(g,L)=L^{k}[f((g-g_{c})L^{1/\nu})+bL^{-1}]$.
In the following section, by expanding the scaling function $f(\Delta gL^{1/\nu})$ (or $f^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)d O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
Ising transition.— We consider the bilayer Ising-Heisenberg model with (2+1) D Ising transition  [37]. The Hamiltonian reads:
$H=-J_{z}\sum_{\alpha=1,2}\sum_{\langle ij\rangle}S_{i,\alpha}^{z}S_{j,\alpha}^{z}+J\sum_{i}{\mathbf{S}}_{i,1}\cdot{\mathbf{S}}_{i,2},$ (7)
in which $\langle ij\rangle$ denotes the nearest-neighbor interaction, and $\alpha=1,2$ is the layer index. $-J_{z}<0$ represents the intralayer ferromagnetic (FM) Ising interaction, while $J>0$ represents the interlayer antiferromagnetic (AF) Heisenberg interaction. For $J/J_{z}\ll 1$, the ground state exhibits Ising order, breaking the $\mathbb{Z}_{2}$ symmetry. For $J/J_{z}\gg 1$, the ground state is in a spin-singlet states. The QCP separating two phases is at $J/J_{z}=3.045(2)$ and belongs to the (2+1)D Ising universality class [37].
We set $J=3.045$ and use $J_{z}$ as the tuning parameter. We then evaluate the 2nd-order RTE and its derivative in the vinicity of the QCP at $J_{z,c}=1$ for various lattice sizes. The results are shown in Fig. 1 (a) and (c). The RTEs of different sizes intersect at QCP, and their derivatives have a peak at QCP. We fit the scaling forms given by Eq. (5) and (6) to the data of RTE and its derivative separately. For the RTE data, we find $J_{z,c}=0.9999(1)$ and $\nu=0.62(1)$, while for the derivative of the DRTE data, we find $J_{z,c}=1.000(3)$ and $\nu=0.63(4)$. These results are consistent with the QCP $J_{z,c}=1$  [37] and the critical exponent $\nu=0.63012$ in the (2+1)D Ising universality class [38, 39, 37]. The universal scaling functions $f(\Delta gL^{1/\nu})$ and $f^{\prime}(\Delta gL^{1/\nu})$ obtained from the data collapse analysis are plotted in Fig. 1 (b) and (d).
Continuous symmetry breaking transitions.— We then investigate QCPs of the dimerized AF Heisenberg model [40, 41, 42]. The Hamiltonian reads:
$H=J_{1}\sum_{\langle ij\rangle}D_{ij}+J_{2}\sum_{\langle ij\rangle^{\prime}}D_{ij},$ (8)
in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction and $\Delta$ is anisotropy parameter. In the regime where $J_{2}/J_{1}\simeq 1$, the ground state exhibits AF order. For $10<\Delta<$ 1, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry. For $\Delta>1$, the order parameter aligns along the $z$-axis, breaking the $\mathbb{Z}_{2}$ spin flip symmetry. When $\Delta=1$, the interaction is isotropic, and the AF order breaks the O(3) symmetry. In the limit $J_{2}/J_{1}\gg 1$, the ground state is the gapped dimer phase. The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $10<\Delta<$ 1/ $\Delta>1$ / $\Delta=1$  [40, 41, 42].
The RTE and its derivative are evaluated around the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality). The results are shown in Fig. 1 (e,g) and (i,k), respectively. In each case, we fix $J_{2}$ and tune $J_{1}$ across the QCP, as described in the figure caption. Similar to the Ising transition case, we find that the RTE values for different lattice sizes intersect at the QCPs, and their derivatives exhibit peaks at the QCPs.
We fit the scaling forms given by Eq. (5) and (6) near the (2+1)D O(2) and O(3) QCPs separately. Around the O(2) QCP, we find $J_{2,c}=0.997(1)$ and $\nu=0.67(2)$ for the RTE data, and $J_{2,c}=0.999(1)$ and $\nu=0.66(2)$ for the DRTE data. These results are consistent with the known QCP and critical exponent: $J_{1,c}=1$ and $\nu=0.6703$ [38, 42].
Near the O(3) QCP, we find $J_{2,c}=0.523(1)$ and $\nu=0.70(1)$ for the RTE data, and $J_{2,c}=0.525(1)$ and $\nu=0.71(2)$ for the DRTE data. These results are also consistent with the known QCP and the critical exponent: $J_{1,c}=0.52337$ and $\nu=0.7073$ [38, 40, 41]. The universal scaling functions near O(2) and O(3) QCPs obtained from the data collapse analysis are shown in Fig. 1 (f), (h) and (g), (m) .
Figure 2: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$, and (b) and (c) the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model. (d–f) Data for the $J-Q_{3}$ model. (g–i) Data for the $J-Q_{2}$ model.
First-order and weakly first-order transitions.— Having established the universal scaling form for (2+1)D QCPs, we then investigate first-order transitions, especially the weakly first-order phase transition presenting significant challenges for traditional local physical quantities.
We first study the 2D checkerboard $J$-$Q$ model, which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [43]. The DRTE around this transition is plotted in Fig. 2 (a). A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side of the transition as $L\to\infty$. This behavior arises because DRTE isolates the singular contribution to the logarithmic partition function (as mentioned above). For comparison, we also calculate the derivative of logarithmic partition functions, $2\partial\ln Z(\beta)/\partial J$ and $\partial\ln Z(2\beta)/\partial J$, as shown in Fig. 2 (b) and (c). The curve seems much more continuous compared with DRTE, because the derivative of logarithmic partition function includes a much larger term obeying volume-law. The phase transition of this model is obviously first-order according to the previous works [43, 23], thus we see that the cuves in (b) and (c) have very small bends around the phase transition point, especially for larger sizes.
The phase transition between AF Néel order and valence-bond solid in 2D quantum magnets can be (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [44, 45]. $J-Q_{2}$ and $J-Q_{3}$ models can realize the DQC in 2D quantum spin systems [46, 47]. This model has no sign problem and is, therefore, amenable to QMC simulations. Numerous QMC studies have shown that the phase transitions of the $J-Q$ model on the 2D lattice are continuous and exhibit emergent SO(5) symmetry, which is a hallmark of DQC [46, 47, 48, 49, 50, 51]. However, the exponents of the $J-Q$ models do not satisfy the conformal bootstrap bounds of a $SO(5)$ critical point [52]. Recent studies have provided growing evidence suggesting that this DQC point is in fact a weakly first-order phase transition [53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. In the following, we demonstrate that DRTE can clearly and simply prove that the phase transition in the $J-Q_{2}$ and $J-Q_{3}$ systems is a first-order phase transition.
The transition points of the two-dimensional square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [46, 47, 63, 62]. The numerical results of the DRTE for these two models are presented in Fig. 2 (d) and (g), where we also show the results for the derivative of logarithmic partition function itself in (e) (f) and (h) (i). The presence of a dominant background term makes it difficult for the derivative of the logarithmic partition function to directly reveal the weakly first-order nature of the phase transition. As shown in Fig. 2 (e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model, the curves appear smooth across the transition point, even for large system sizes. On the other hand, the DRTEs of both the $J-Q_{3}$ and $J-Q_{2}$ models exhibit clear discontinuities near the phase transition points. The results, by definition, prove that the phase transitions in both the $J-Q_{2}$ and $J-Q_{3}$ models are (weakly) first-order phase transitions. Notably, compared to $J-Q_{3}$ model, $J-Q_{2}$ model exhibits stronger finite-size effects, making it more difficult to determine its first-order nature prior to our study [62]. Our findings demonstrate that even weakly first-order phase transitions can be clearly detected using DRTE, even in normal sizes.
Conclusion and discussions.— We have proposed that the RTE only contains the singular term of logarithmic partition function, which is highly sensitive around phase transition points. Then we have systematically studied the RTE and its derivative around the QCPs in 2D quantum spin systems. We derived universal scaling forms for the RTE and DRTE, which capture the numerical results and yield the critical exponent $\nu$ through data collapse analysis near (2+1)D QCPs. More importantly, by definition, DRTE can clearly detect weak first-order phase transitions associated with deconfined quantum criticality, a task that is highly challenging for traditional local physical quantities.
Notably, RTE and DRTE is a universally applicable quantity because its definition does not depend on specific models or the precise definition of order parameters. Studying them has the potential to characterize phase transitions that are challenging for traditional local physical quantities, such as topological phase transitions, and to discover new universal critical exponents. Our work introduces a new and powerful paradigm for characterizing quantum phase transitions, including continuous phase transitions, first-order phase transitions, and the particularly challenging weakly first-order phase transitions.
Acknowledgements.
We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang. Z.W. thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
Supplemental Materials
Appendix A Method
The $n$th order Rényi entropy is defined as $S^{(n)}=\frac{1}{1-n}\ln\mathrm{Tr}(\rho^{n})$ (or $\frac{1}{1-n}\ln\mathrm{Tr}(\rho_{A}^{n})$), where $\rho=e^{-\beta H}/Z$ and $Z=\mathrm{Tr}e^{-\beta H}$ (with $H$ being the Hamiltonian ). Here, $\rho_{A}=\mathrm{Tr}_{B}\rho$ represents the reduced density matrix of subsystem $A$ coupled with an environment $B$. The Rényi entanglement entropy is defined using $\rho_{A}$, while the Rényi thermal entropy (RTE) is based on $\rho$. In this work, we focus on RTE. From the above definitions, it is evident that the R’enyi entropy can be expressed as the ratio of partition functions at two different temperatures: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}$ (for simplicity, we consider $S^{(2)}$ in this work).
The partition functions can be calculated using the bipartite reweight-annealing algorithm [23, 27, 64, 65] in conjunction with stochastic series expansion (SSE) QMC simulations [30, 12, 31, 32, 34, 33]. The key idea is to transform the problem of solving the ratio of partition functions with different parameter into the problem of reweighting:
$\frac{Z(J^{\prime})}{Z(J)}=\bigg{\langle}\frac{W(J^{\prime})}{W(J)}\bigg{\rangle}_{Z(\beta,J)orZ(n\beta,J)}$ (S1)
where the $\langle...\rangle_{Z(\beta,J)orZ(n\beta,J)}$ indicates that the QMC samplings is performed under the manifold $Z(\beta,J)orZ(n\beta,J)$ at parameter $J$ (note parameter $J$ is physical parameter here). $W(J^{\prime})$ and $W(J)$ represent the weights of a same configuration but under different parameters $J^{\prime}$ and $J$.
In this frame, we can designed the incremental process along the path composed of real physical parameters  [^1] It was virtual path before. In other words, all intermediates are physical partition functions at corresponding parameters. The continuously incremental trick here are introduce as follows:
$\frac{Z(J^{\prime})}{Z(J)}=\prod_{i=0}^{N-1}\frac{Z(J_{i+1})}{Z(J_{i})}$ (S2)
where $J_{N}=J^{\prime}$ and $J_{0}=J$, others $J_{i}$ locate between the two in sequence. In realistic simulation, we are able to gain any ratio ${Z(J^{\prime})}/{Z(J)}$ in this way. But $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can not be obtained. The antidote comes from some known reference point $Z(n\beta,J)/Z^{n}(\beta,J)$. Considering a zero-temperature system and assuming the degeneracy of the ground state is $de$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln de$+$c((J-J_{c})L^{(1/\nu)})$, where $c$ represents the critical corrections, $J_{c}$ represents the critical point, and $\nu$ is the critical exponent of the correlation length. Therefore, we can use some states with known degeneracy as reference points, such as product states and discrete symmetry-breaking states.
The RTE derivative of $J$ is
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}\frac{dZ(n\beta)/dJ}{Z(n\beta)}-n\frac{dZ(\beta)/dJ}{Z(\beta)}\bigg{]}$ (S3)
According to $Z(\beta)=\mathrm{Tr}e^{-\beta H}$, we have
$\frac{dZ}{dJ}=\mathrm{Tr}[-\beta\frac{dH}{dJ}e^{-\beta H}]$ (S4)
Thus
$\frac{dZ/dJ}{Z}=-\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}$ (S5)
Similarly, we can obtain
$\begin{split}\frac{dZ(n\beta))/dJ}{Z(n\beta))}&=-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta))}\end{split}$ (S6)
Therefore, the RTE derivative can be rewritten as
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta))}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}\bigg{]}$ (S7)
The equations above is general and doesn’t depend on detailed quantum Monte Carlo (QMC) methods. Then let us discuss how to calculate them in stochastic series expansion (SSE) simulation. For convenience, we fix the Rényi index $n=2$ and choose the dimerized Heisenberg model in main text as the example for explaining technical details. The Hamiltonian is
$H=J_{1}\sum_{\langle ij\rangle}S_{i}S_{j}+J_{2}\sum_{\langle ij\rangle}S_{i}S_{j}$ (S8)
In the following, we fix $J_{2}=1$ and leave $J_{1}$ as the tunable parameter. Note the Hamiltonian is a linear function of $J_{1}$, that means $dH/dJ_{1}=H_{J_{1}}/J_{1}$ in which $H_{J_{1}}$ is the $J_{1}$ term in Hamiltonian. Then the RTE derivative can be simplified as
$\frac{dS^{(2)}}{dJ_{1}}=\bigg{[}2\beta\bigg{\langle}\frac{H_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z(n\beta)}-2\beta\bigg{\langle}\frac{H_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z}\bigg{]}$ (S9)
In the SSE frame, it is easy to obtain $\langle H\rangle=\langle-n_{op}/\beta\rangle$  [^2] How to measure the energy in SSE has been carefully explained in Prof. Sandvik’s tutorial http://physics.bu.edu/~sandvik/programs/ssebasic/ssebasic.html, where $n_{op}$ is the number of the concerned Hamiltonian operators. Thus the Eq. (S9) can be further simplified to
$\frac{dS^{(2)}}{dJ_{1}}=\bigg{[}-\bigg{\langle}\frac{n_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z(n\beta)}+2\bigg{\langle}\frac{n_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z}\bigg{]}$ (S10)
where $n_{J_{1}}$ means the number of $J_{1}$ operators including both diagonal and off-diagonal ones. It’s worth noting that there is no “$2$” anymore in the $\langle...\rangle_{Z(n\beta)}$ term, because $n_{J_{1}}$ here contains two replicas’ operators which has already been doubled actually.

### v2

@@
-Extracting the singularity of the logarithmic partition function
+Extracting the singularity of a logarithmic partition function to investigate quantum phase transitions
 Abstract
-In principle, the logarithm of the partition function and its derivatives can be used to rigorously determine the nature of phase transitions.
-However, in practice, this approach is often ineffective in numerical calculations, since resolving the singularities of the logarithmic partition function requires extremely high-precision data—especially when distinguishing between weakly first-order phase transitions and deconfined quantum critical points.
-Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the leading volume-law term.
+In principle, the logarithm of a partition function and its derivatives can be used to rigorously determine the nature of phase transitions.
+However, this approach is often ineffective in numerical calculations in practice, since resolving the singularity of a logarithmic partition function requires extremely high-precision data—especially when distinguishing weakly first-order phase transitions and deconfined quantum critical points.
+Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the analytic terms.
 This is achieved by computing the Rényi thermal entropy (RTE) and its derivative (DRTE).
 We have derived the scaling relations for RTE and DRTE, which successfully determine the phase transition points and critical exponents through data collapse in various quantum many-body systems.
-Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in $J\text{-}Q$ models, which were initially thought to be continuous deconfined quantum phase transitions.
+Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in JQJ-Q models, which were initially thought to be continuous deconfined quantum phase transitions.
 The discontinuity of these transitions is challenging to detect using traditional local physical quantities.
 This work provides a new and powerful paradigm for studying quantum many-body phase transitions.
-Introduction.— Quantum phase transitions, spontaneous symmetry breaking, and the universality classes of critical phenomena are central concepts in condensed matter and statistical physics [1, 2, 3].
-The type of phase transition is well defined by the singularity of free energy and its derivative.
-For a first-order phase transition, the first-order derivative of free energy is discontinuous.
-Otherwise, if its higher-order derivative is discontinuous, it is named a continuous phase transition.
-In principle, if the partition function can be calculated exactly, the phase transition will be fixed strictly according to the singularity.
-However, in practical numerical simulations, directly detecting such singularities through the free energy and its derivatives is difficult, as it requires extremely high-precision data.
-This challenge is particularly pronounced when distinguishing between weakly first-order phase transitions and deconfined quantum critical points  [4, 5, 6, 7, 8, 9, 10].
-A major reason for this difficulty is that the singular contribution is typically overshadowed by a much larger, non-singular term that scales with the system volume.
-How to extract the singularity remains a significant challenge in theoretically computational physics.
-Considering a quantum phase transition ($\beta\rightarrow\infty$), the singularity of free energy can be replaced by focusing on the logarithmic partition function.
-According to the finite-size scaling hypothesis [11, 12], an observable which is singular at $g_{c}$ (the parameter $g$ of Hamiltonian at the phase transition point) scales with the system size close to $g_{c}$ as a power of $L$ multiplied by a non-singular function of the ratio $\xi/L$, where $\xi$ is the correlation length and $L$ is the system size.
-Any singular quantity (not necessarily divergent) should thus be of the form:
-$Q(\Delta g,L)=L^{\sigma}f(\Delta gL^{1/\nu})$ (1)
-where $\Delta g=g-g_{c}$, $\nu$ is the critical exponent of the correlation length and $\sigma$ is an undetermined power.
-Since the singular part of the free energy is scale-invariant ($\sigma=0$[13, 11], the logarithmic partition function can be written in the following form:
-$\ln Z(g)=a(g)L^{d}+f(\Delta gL^{1/\nu})$ (2)
-where $Z$ is a partition function, $d$ is the spatial dimension and $a(g)L^{d}$ represents the leading term obeying a volume law.
-Here, we see that this dominant volume law term often masks the singular contribution of $f$, which explains the difficulty in numerically detecting the phase transition through such singularities.
-An approach to extract the singular term $f$ is to construct a quantity in which the volume-law contributions cancel out, leaving only the singular part.
-Inspired by the method used to calculate topological entanglement entropy [14, 15], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\rightarrow\infty$) isolates the singular term $f$, with the dominant volume-law contributions canceled out.
+Introduction.— Quantum phase transitions, associated spontaneous symmetry breaking, and universality classes are central concepts in condensed matter and statistical physics [1, 2, 3].
+The nature of a phase transition is determined by the singularity in the free energy or its derivatives: a first-order transition is marked by a discontinuity in the first derivative, while a continuous transition involves singular higher-order derivatives.
+In principle, an exact calculation of the partition function would allow one to identify the transition based on these singularities.
+However, in practice, detecting such features in numerical simulations is challenging, as it requires extremely high-precision data.
+This issue is particularly severe when distinguishing weakly first-order transitions from deconfined quantum critical points [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
+A key difficulty arises because the singular contribution is often overwhelmed by a much larger, non-singular term that scales with the system volume.
+Extracting the singular part therefore remains a major challenge in computational physics.
+For a quantum phase transition (inverse temperature $\beta\rightarrow\infty$; in the realistic simulations, the finite-size system gap $\Delta_{L}\sim\frac{1}{L^{z}}$, $\beta\sim L^{z}$ is usually used for probing the ground state properties, in which $L$ is the system size and $z$ is the dynamical critical exponent [17] ), the singularity of the free energy can be equivalently studied through the logarithm of the partition function.
+According to the finite-size scaling hypothesis [18, 17], a singular observable $Q$ (not necessarily divergent) near the critical point $g_{c}$ scales with system size as:
+$Q(\Delta g,L)=L^{\sigma}\tilde{S}(\Delta gL^{1/\nu})$ (1)
+where $\Delta g=g-g_{c}$, $\nu$ is is the correlation length critical exponent, $\sigma$ is a scaling power and $\tilde{S}$ is a universal scaling function.
+Since the singular part of the free energy is scale-invariant ($\sigma=0$[19, 18], the logarithm of the partition function can be expressed as:
+$\ln Z(g,\beta,L)=a(g)\beta L^{d}+\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}(\Delta gL^{1/\nu})$ (2)
+where $Z$ is a partition function, $d$ is the spatial dimension, $L$ is the system length, inverse temperature $\beta\sim L^{z}$ and $a(g)\beta L^{d}$ represents the leading term obeying a volume law, and $\mathcal{A}_{sub}(g,L,\beta)=\beta(a_{1}(g)L^{d-1}+a_{2}(g)L^{d-2}+a_{3}(g)L^{d-3}+\cdots)$ accounts for sub-leading analytic contributions ($a_{i}$ can be zero or non-zero).
+The dominant volume-law term often obscures the singular contribution $\tilde{S}$, which explains the difficulty in detecting phase transitions numerically through such singularities.
+To isolate the singular term $\tilde{S}$, one can construct a quantity in which the dominant analytic contributions cancel out.
+Inspired by the method used to extract topological entanglement entropy [20, 21], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\sim L^{z}$) can serve this role.
 Specifically, we argue that Eq.
-(2) can be rewritten as $\ln Z(2\beta)=2a(g)L^{d}+f$ and $2\ln Z(\beta)=2a(g)L^{d}+2f$.
-Coincidentally, we find that this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [16, 17, 18, 19, 20, 21, 22]:
+(2) can be rewritten as $\ln Z(2\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}$ and $2\ln Z(\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+2\tilde{S}$.
+Coincidentally, this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [22, 23, 24, 25, 26, 27, 28]:
 $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$ (3)
-This definition can be generalized to an arbitrary order $n$, leading to the $n$-th order RTE: $S^{(n)}=\frac{1}{1-n}\ln[Z(n\beta)/Z(\beta)^{n}]$.
-For simplicity, we focus on the 2nd-order RTE ( i.e., $n=2$) throughout this paper.
-In this work, assisted by our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [23, 24, 25, 26, 27, 28, 29], we demonstrate the validity of our conjecture above: the RTE is indeed a dimensionless quantity that captures only the singular contribution.
-Furthermore, we observe a more pronounced singularity in the derivative of the RTE (DRTE) near the quantum critical point (QCP).
-We propose universal scaling relations for both the RTE and DRTE, which are confirmed by data collapse analyses performed near (2+1)D QCPs belonging to the O($N$) universality classes, with $1\leq N\leq 3$.
-More interestingly, since the leading non-singular term of $\ln Z$ is canceled out in the DRTE, it becomes a particularly powerful tool for detecting weakly first-order phase transitions even in normal sizes, which are notoriously difficult to resolve using conventional numerical approaches.
+This definition generalizes naturally to arbitrary order $n$ : $S^{(n)}=\frac{1}{1-n}\ln[Z(n\beta)/Z(\beta)^{n}]$.
+For clarity and simplicity, we focus on the second-order RTE ($n=2$) throughout this work.
+In this work, using our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [29, 30, 31, 32, 33, 34, 35], we validate the conjecture that the Rényi thermal entropy (RTE) is a dimensionless quantity capturing only the singular contribution.
+Moreover, its derivative (DRTE) exhibits a more pronounced singularity near the quantum critical point (QCP).
+We propose universal scaling forms for both RTE and DRTE, which are confirmed through data collapse analyses at (2+1)-dimensional (D) QCPs in the O($N$) universality classes, with $1\leq N\leq 3$.
+Notably, because the leading non-singular term in $\ln Z$ is canceled in DRTE, it provides a powerful diagnostic for weakly first-order phase transitions, even in relatively small systems (up to $L=56$), which are significant challenges to resolve using conventional numerical approaches.
 Models and method.— We investigate quantum spin models with continuous QCPs in the (2+1)D O($N$) universality classes, where $1\leq N\leq 3$.
-Additionally, we explore first-order transitions, with a particular focus on the challenging weakly first-order phase transitions (e.g., some controversial deconfined quantum critical points).
-Detailed descriptions of the model Hamiltonians will be provided in the following sections.
-As shown by the definitions above, the 2nd-order Rényi thermal entropy can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$.
-The partition functions can be calculated using the bipartite reweight-annealing algorithm [23, 24] in conjunction with stochastic series expansion (SSE) QMC simulations [30, 12, 31, 32, 33, 34, 35].
-Furthermore, following a procedure similar to that used in calculating the derivative of Rényi entanglement entropy and negativity [23, 36], the derivative of the RTE with respect to a general parameter $J$ can be written as:
-$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
-In this expression, the first average is evaluated on the ensemble corresponding to $Z(n\beta)$, while the second is based on $Z(\beta)$.
-In our simulations, the inverse temperature $\beta$ is set to be linearly proportional to the lattice size $L$ (i.e., $\beta=L$) in order to approach the quantum critical regime.
-Further methodological details are provided in the Supplementary Materials (SM).
-Finite-size scaling form.— Considering a zero-temperature system and assuming the ground-state degeneracy is $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$+$f(\Delta gL^{1/\nu})$, where $f$ represents the critical corrections.
-As shown in Fig.1, $S^{(2)}$ can be used to probe both the degeneracy of discrete symmetry-breaking and the tower of excited states (TOS) associated with continuous symmetry-breaking.
+We also examine first-order transitions, with a focus on weakly first-order cases, including some debated deconfined quantum critical points.
+The model Hamiltonians are described in detail in the following sections.
+The second-order RTE can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$.
+We compute the partition functions using the bipartite reweight-annealing algorithm [29, 30] in combination with stochastic series expansion (SSE) QMC simulations [36, 17, 37, 38, 39, 40, 41].
+Following a method analogous to that used for the derivative of Rényi entanglement entropy and negativity [29, 42], the derivative of the RTE with respect to a general parameter $J$ is given by:
+$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
+Here, the two expectation values are evaluated separately using ensembles corresponding to $Z(n\beta)$ and $Z(\beta)$, respectively.
+While the RTE is calculated using the reweight-annealing QMC method, the derivative (DRTE) is directly obtained from the two expectation values in Eq.
+(4); see Supplementary Materials (SM) for details.
+Considering the dynamical critical exponents $z=1$ of the models studied in this work, the $\beta=L$ is set to approach the quantum critical regime.
+Finite-size scaling form.—For a zero-temperature system with ground-state degeneracy $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$$\tilde{S}(\Delta gL^{1/\nu})$, where $\tilde{S}$ represents the critical corrections.
+As shown in Fig.1, $S^{(2)}$ effectively probes both the degeneracy of discrete symmetry-breaking phases and the tower of excited states (TOS) associated with continuous symmetry-breaking.
+Near criticality, the logarithmic partition function consists of a dominant non-singular term $a(J)\beta L^{d}$ and a dimensionless singular term $\tilde{S}(\Delta gL^{1/\nu})$.
+The leading non-singular term obscures the critical behavior, making it difficult to extract the singular part directly from the partition function.
 More interestingly, as shown in Fig.
-1, we observe that the curves for different system sizes cross at the critical point, which implies that $S^{(2)}$ is indeed a dimensionless quantity, consistent with our argument in the introduction.
-Near the critical point, the logarithmic partition function can be decomposed into a non-singular part $a(J)L^{d}$ and a dimensionless singular term $f(\Delta gL^{1/\nu})$.
-The non-singular contribution is the leading term, which makes it difficult to extract critical information directly from the partition function.
-As shown in Fig.
-1, the crossings of curves for different sizes indicate that the RTE is independent of the system size.
-This provides strong numerical support for our argument that the leading volume-law contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel out in the definition of RTE, leaving only the dimensionless singular part.
-Additionally, we observe that the DRTE exhibits a stronger divergence near the QCP.
-According to the finite-size scaling hypothesis  [11, 12], any singular quantity near a critical point should obey a scaling form as given in Eq.(1).
-In particular, the dimensionless RTE ($\sigma=0$), is expected to follow the scaling form:
-$S^{(n)}(\Delta g,L)=f(\Delta gL^{1/\nu}).$ (5)
+1, the curves of $S^{(2)}$ for different system sizes intersect at the critical point, indicates that the RTE is independent of system size, providing strong numerical evidence that the leading non-singular contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel exactly in the definition of RTE.
+This cancellation isolates the dimensionless singular term.
+Moreover, the DRTE displays a more pronounced divergence near the QCPs, further highlighting its sensitivity to critical fluctuations.
+According to the finite-size scaling hypothesis  [18, 17], any singular quantity near a critical point follows a scaling form as given in Eq.(1).
+In particular, the dimensionless nature of the RTE ($\sigma=0$), implies the scaling form:
+$S^{(n)}(\Delta g,L)=\tilde{S}(\Delta gL^{1/\nu})$ (5)
 From Eq. (5), we can readily derive the scaling behavior of the DRTE:
-$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=f^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
-where $f^{\prime}(x)$ is $\partial f(x)/\partial x$ and $x$ is a general parameter.
-Figure 1: (a) The 2nd-order RTE and (c) its derivative, and (b) and (d) the data collapse analysis of the RTE and DRTE near the QCP of the bilayer Ising-Heisenberg model (with (2+1)D Ising universality class).
+$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
+where $\tilde{S}^{\prime}(x)$ is $\partial\tilde{S}(x)/\partial x$ and $x$ is a general parameter.
+Figure 1: (a,c) Second-order RTE and its derivative DRTE, and (b,d) corresponding data collapse analysis near the QCP of the bilayer Ising–Heisenberg model [(2+1)D Ising universality].
 We set $J=3.045$ and take $J_{z}$ as the tuning parameter.
-(e–h) Data of the dimerized anisotropic Heisenberg model with $\Delta=0.9$ (with (2+1)D O(2) universality class).
-We set $J_{2}=2.1035$, and take $J_{1}$ as the tuning parameter.
-(i–m) Data of the dimerized isotropic Heisenberg model with $\Delta=1$ (with (2+1)D O(3) universality class).
-We set $J_{2}=1$, and take $J_{1}$ as the tuning parameter.
-To achieve good data collapse, a subleading correction is also included: $Q(g,L)=L^{k}[f((g-g_{c})L^{1/\nu})+bL^{-1}]$.
-In the following section, by expanding the scaling function $f(\Delta gL^{1/\nu})$ (or $f^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)d O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
-Ising transition.— We consider the bilayer Ising-Heisenberg model with (2+1) D Ising transition  [37]. The Hamiltonian reads:
+(e–h) Results for the dimerized anisotropic Heisenberg model with $\Delta=0.9$ [(2+1)D O(2) universality], using $J_{2}=2.1035$ and tuning $J_{1}$.
+(i–m) Results for the dimerized isotropic Heisenberg model with $\Delta=1$ [(2+1)D O(3) universality], using $J_{2}=1$ and tuning $J_{1}$.
+To achieve good data collapse, we include a subleading correction: $Q(g,L)=L^{k}[\tilde{S}((g-g_{c})L^{1/\nu})+bL^{-\omega}]$.
+$\omega$ is close to 1, and to reduce the numerical instability caused by increasing the number of fitting parameters, we set $\omega=1$.
+For $Q=S^{(2)}$, $k=0$.
+$k=1/\nu$ when $Q$ is DRTE.
+In the following section, by expanding the scaling function $\tilde{S}(\Delta gL^{1/\nu})$ (or $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)D O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
+Ising transition.— We study the bilayer Ising-Heisenberg model with a (2+1)D Ising QCP [43]. The Hamiltonian is given by:
 $H=-J_{z}\sum_{\alpha=1,2}\sum_{\langle ij\rangle}S_{i,\alpha}^{z}S_{j,\alpha}^{z}+J\sum_{i}{\mathbf{S}}_{i,1}\cdot{\mathbf{S}}_{i,2},$ (7)
-in which $\langle ij\rangle$ denotes the nearest-neighbor interaction, and $\alpha=1,2$ is the layer index.
-$-J_{z}<0$ represents the intralayer ferromagnetic (FM) Ising interaction, while $J>0$ represents the interlayer antiferromagnetic (AF) Heisenberg interaction.
-For $J/J_{z}\ll 1$, the ground state exhibits Ising order, breaking the $\mathbb{Z}_{2}$ symmetry.
-For $J/J_{z}\gg 1$, the ground state is in a spin-singlet states.
+where $\langle ij\rangle$ denotes the nearest-neighbor interaction, and $\alpha=1,2$ labels the two layers.
+$J_{z}<0$ represents the intralayer ferromagnetic (FM) Ising interaction, while $J>0$ represents the interlayer antiferromagnetic (AF) Heisenberg interaction.
+For $J/J_{z}\ll 1$, the ground state exhibits $\mathbb{Z}_{2}$ symmetry-breaking Ising order.
+In the opposite limit, $J/J_{z}\gg 1$, the system forms interlayer spin-singlet states.
 The QCP separating two phases is at $J/J_{z}=3.045(2)$ and belongs to the (2+1)D Ising universality class [43].
-We set $J=3.045$ and use $J_{z}$ as the tuning parameter.
-We then evaluate the 2nd-order RTE and its derivative in the vinicity of the QCP at $J_{z,c}=1$ for various lattice sizes.
-The results are shown in Fig.
-1 (a) and (c).
-The RTEs of different sizes intersect at QCP, and their derivatives have a peak at QCP.
-We fit the scaling forms given by Eq.
-(5) and (6) to the data of RTE and its derivative separately.
-For the RTE data, we find $J_{z,c}=0.9999(1)$ and $\nu=0.62(1)$, while for the derivative of the DRTE data, we find $J_{z,c}=1.000(3)$ and $\nu=0.63(4)$.
-These results are consistent with the QCP $J_{z,c}=1$  [37] and the critical exponent $\nu=0.63012$ in the (2+1)D Ising universality class [38, 39, 37].
-The universal scaling functions $f(\Delta gL^{1/\nu})$ and $f^{\prime}(\Delta gL^{1/\nu})$ obtained from the data collapse analysis are plotted in Fig.
-1 (b) and (d).
+We set $J=3.045$ and use $J_{z}$ as the tuning parameter, studying the second-order RTE and DRTE near the QCP at $J_{z,c}=1$ for various lattice sizes.
+The results are shown in Figs.
+1(a) and (c).
+The RTE curves for different system sizes intersect at the QCP, while the DRTE exhibits a clear peak at the critical point.
+We fit the scaling forms in Eqs.
+(5) and (6) to the RTE and DRTE data, respectively.
+The fits yield $J_{z,c}=0.9999(1)$, $\nu=0.62(1)$ from RTE, and $J_{z,c}=1.000(3)$, $\nu=0.63(4)$ from DRTE.
+These values are agreement with the known QCP $J_{z,c}=1$  [43] and the (2+1)D Ising universality class value $\nu=0.63012$  [44, 45, 43].
+The resulting universal scaling functions $\tilde{S}(\Delta gL^{1/\nu})$ and $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$, obtained from data collapse, are displayed in Figs.
+1(b) and (d), respectively.
 Continuous symmetry breaking transitions.— We then investigate QCPs of the dimerized AF Heisenberg model [46, 47, 48]. The Hamiltonian reads:
 $H=J_{1}\sum_{\langle ij\rangle}D_{ij}+J_{2}\sum_{\langle ij\rangle^{\prime}}D_{ij},$ (8)
-in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction and $\Delta$ is anisotropy parameter.
+in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction with anisotropy parameter $\Delta$.
 In the regime where $J_{2}/J_{1}\simeq 1$, the ground state exhibits AF order.
-For $10<\Delta<$ 1, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry.
+For $0<\Delta<1$, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry.
 For $\Delta>1$, the order parameter aligns along the $z$-axis, breaking the $\mathbb{Z}_{2}$ spin flip symmetry.
@@
 In the limit $J_{2}/J_{1}\gg 1$, the ground state is the gapped dimer phase.
-The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $10<\Delta<$ 1/ $\Delta>1$ / $\Delta=1$  [40, 41, 42].
-The RTE and its derivative are evaluated around the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality).
-The results are shown in Fig.
-1 (e,g) and (i,k), respectively.
-In each case, we fix $J_{2}$ and tune $J_{1}$ across the QCP, as described in the figure caption.
-Similar to the Ising transition case, we find that the RTE values for different lattice sizes intersect at the QCPs, and their derivatives exhibit peaks at the QCPs.
-We fit the scaling forms given by Eq.
-(5) and (6) near the (2+1)D O(2) and O(3) QCPs separately.
-Around the O(2) QCP, we find $J_{2,c}=0.997(1)$ and $\nu=0.67(2)$ for the RTE data, and $J_{2,c}=0.999(1)$ and $\nu=0.66(2)$ for the DRTE data.
-These results are consistent with the known QCP and critical exponent: $J_{1,c}=1$ and $\nu=0.6703$ [38, 42].
-Near the O(3) QCP, we find $J_{2,c}=0.523(1)$ and $\nu=0.70(1)$ for the RTE data, and $J_{2,c}=0.525(1)$ and $\nu=0.71(2)$ for the DRTE data.
-These results are also consistent with the known QCP and the critical exponent: $J_{1,c}=0.52337$ and $\nu=0.7073$ [38, 40, 41].
-The universal scaling functions near O(2) and O(3) QCPs obtained from the data collapse analysis are shown in Fig.
-1 (f), (h) and (g), (m) .
+The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $0<\Delta<1$/ $\Delta>1$ / $\Delta=1$  [46, 47, 48].
+We evaluate the RTE and its derivative near the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality), with results shown in Figs.
+1(e,g) and (i,k), respectively.
+In each case, $J_{2}$ is fixed and $J_{1}$ is tuned across the transition, as detailed in the caption.
+Similar to the Ising case, the RTE curves for different system sizes intersect at the QCPs, and their derivatives exhibit clear peaks at these points.
 Figure 2: We set $Q=1$ and use $J$ as the tuning parameter.
-(a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$, and (b) and (c) the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model.
+(a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$.
+The inset shows data on the right side of the transition point.
+(b) and (c) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model.
 (d–f) Data for the $J-Q_{3}$ model.
 (g–i) Data for the $J-Q_{2}$ model.
-First-order and weakly first-order transitions.— Having established the universal scaling form for (2+1)D QCPs, we then investigate first-order transitions, especially the weakly first-order phase transition presenting significant challenges for traditional local physical quantities.
-We first study the 2D checkerboard $J$-$Q$ model, which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [43].
+The raw values of the partition function derivative vary significantly across system sizes (see SM), making direct comparison difficult.
+To resolve this, we normalize the data by $L^{3}$, corresponding to $\beta L^{2}=L^{3}$.
+We apply the scaling forms in Eqs.
+(5) and (6) to data near the (2+1)D O(2) and O(3) QCPs.
+For the O(2) QCP, fitting the RTE yields $J_{1,c}=0.997(1)$, $\nu=0.67(2)$, while the DRTE gives $J_{1,c}=0.999(1)$, $\nu=0.66(2)$.
+These values are consistent with the known critical parameters $J_{1,c}=1$ and $\nu=0.6703$ [44, 48].
+For the O(3) QCP, fitting the RTE data yields $J_{1,c}=0.523(1)$, $\nu=0.70(1)$, while the DRTE gives $J_{1,c}=0.525(1)$, $\nu=0.71(2)$.
+These values are consistent with the known critical parameters $J_{1,c}=0.52337$ and $\nu=0.7073$ [44, 46, 47].
+The universal scaling functions near the O(2) and O(3) QCPs, obtained from data collapse, are shown in Figs.
+1(f,h) and (j,m), respectively.
+First-order and weakly first-order transitions.— Having established the universal scaling behavior at (2+1)D QCPs, we now turn to first-order phase transitions, with a focus on weakly first-order transitions, which presents significant challenges for traditional local physical quantities.
+We first study the 2D checkerboard $J$-$Q$ model ( see SM for model details), which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [49].
 The DRTE around this transition is plotted in Fig.
 2 (a).
-A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side of the transition as $L\to\infty$.
+A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side as $L\to\infty$.
 This behavior arises because DRTE isolates the singular contribution to the logarithmic partition function (as mentioned above).
-For comparison, we also calculate the derivative of logarithmic partition functions, $2\partial\ln Z(\beta)/\partial J$ and $\partial\ln Z(2\beta)/\partial J$, as shown in Fig.
+For comparison, we also calculate the derivative of logarithmic partition functions, as shown in Figs.
 2 (b) and (c).
-The curve seems much more continuous compared with DRTE, because the derivative of logarithmic partition function includes a much larger term obeying volume-law.
-The phase transition of this model is obviously first-order according to the previous works [43, 23], thus we see that the cuves in (b) and (c) have very small bends around the phase transition point, especially for larger sizes.
-The phase transition between AF Néel order and valence-bond solid in 2D quantum magnets can be (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [44, 45].
-$J-Q_{2}$ and $J-Q_{3}$ models can realize the DQC in 2D quantum spin systems [46, 47].
-This model has no sign problem and is, therefore, amenable to QMC simulations.
-Numerous QMC studies have shown that the phase transitions of the $J-Q$ model on the 2D lattice are continuous and exhibit emergent SO(5) symmetry, which is a hallmark of DQC [46, 47, 48, 49, 50, 51].
-However, the exponents of the $J-Q$ models do not satisfy the conformal bootstrap bounds of a $SO(5)$ critical point [52].
-Recent studies have provided growing evidence suggesting that this DQC point is in fact a weakly first-order phase transition [53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
-In the following, we demonstrate that DRTE can clearly and simply prove that the phase transition in the $J-Q_{2}$ and $J-Q_{3}$ systems is a first-order phase transition.
-The transition points of the two-dimensional square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [46, 47, 63, 62].
-The numerical results of the DRTE for these two models are presented in Fig.
-2 (d) and (g), where we also show the results for the derivative of logarithmic partition function itself in (e) (f) and (h) (i).
-The presence of a dominant background term makes it difficult for the derivative of the logarithmic partition function to directly reveal the weakly first-order nature of the phase transition.
-As shown in Fig.
-2 (e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model, the curves appear smooth across the transition point, even for large system sizes.
-On the other hand, the DRTEs of both the $J-Q_{3}$ and $J-Q_{2}$ models exhibit clear discontinuities near the phase transition points.
-The results, by definition, prove that the phase transitions in both the $J-Q_{2}$ and $J-Q_{3}$ models are (weakly) first-order phase transitions.
-Notably, compared to $J-Q_{3}$ model, $J-Q_{2}$ model exhibits stronger finite-size effects, making it more difficult to determine its first-order nature prior to our study [62].
-Our findings demonstrate that even weakly first-order phase transitions can be clearly detected using DRTE, even in normal sizes.
-Conclusion and discussions.— We have proposed that the RTE only contains the singular term of logarithmic partition function, which is highly sensitive around phase transition points.
-Then we have systematically studied the RTE and its derivative around the QCPs in 2D quantum spin systems.
-We derived universal scaling forms for the RTE and DRTE, which capture the numerical results and yield the critical exponent $\nu$ through data collapse analysis near (2+1)D QCPs.
-More importantly, by definition, DRTE can clearly detect weak first-order phase transitions associated with deconfined quantum criticality, a task that is highly challenging for traditional local physical quantities.
-Notably, RTE and DRTE is a universally applicable quantity because its definition does not depend on specific models or the precise definition of order parameters.
-Studying them has the potential to characterize phase transitions that are challenging for traditional local physical quantities, such as topological phase transitions, and to discover new universal critical exponents.
-Our work introduces a new and powerful paradigm for characterizing quantum phase transitions, including continuous phase transitions, first-order phase transitions, and the particularly challenging weakly first-order phase transitions.
+These curves appear much smoother across the transition compared with DRTE, due to the dominance of a non-singular, volume-law term.
+The phase transition of this model is obviously first-order according to the previous works [49, 29], thus we see that the curves in (b) and (c) have very small kinks around the phase transition point, especially for larger sizes.
+The phase transition between AF Néel order and valence-bond solid (VBS) phase in 2D quantum magnets can appear (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [4, 5].
+$J-Q_{2}$ and $J-Q_{3}$ models are prototypical examples that realize such transitions in 2D quantum spin systems [50, 51].
+These models are free of the sign problem, making them ideal for QMC simulations.
+Extensive QMC studies have shown that the phase transitions of the JQJ-Q models on the 2D lattice are continuous and exhibit emergent SO(5) symmetry, which is a hallmark of DQC [50, 51, 52, 53, 54, 55].
+However, the critical exponents observed in these models do not satisfy the conformal bootstrap bounds expected for SO(5) critical point [56].
+Recent works have increasingly suggested that this DQC scenario may instead correspond to a weakly first-order transition or as a nonunitary CFT  [7, 11, 57, 58, 12, 59, 60, 61, 8, 15, 62, 63, 64, 65, 16].
+In the following, we show that the DRTE provides a clear and direct signature that the transitions in the $J-Q_{2}$ and $J-Q_{3}$ models ( see SM for models’ details) are indeed first-order, even in rather small systems (up to $L=56$).
+The transition points of the 2D square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [50, 51, 66, 63].
+The DRTE results for both models are shown in Figs.
+2 (d) and (g), along with the derivatives of the logarithmic partition function in panels (e,f) and (h,i).
+Due to the dominant background contribution, the derivative of the logarithmic partition function appears smooth across the transition, even for large system sizes, as seen in Figs.
+2(e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model.
+This shows that the derivative of the logarithmic partition function can not be directly reveal the weakly first-order nature.
+In contrast, the DRTE reveals clear discontinuities at the transition points, providing direct evidence that both transitions are (weakly) first-order.
+Notably, the $J-Q_{2}$ model exhibits stronger finite-size effects than $J-Q_{3}$, which has made its first-order nature more difficult to resolve in previous studies [63].
+Our results demonstrate that DRTE can unambiguously detect even weakly first-order transitions, in relatively small systems (up to $L=56$).
+Conclusion and discussions.— We proposed that the RTE isolates the singular contribution of the logarithmic partition function, making it highly sensitive near phase transitions.
+We systematically studied the RTE and its derivative DRTE around QCPs in 2D quantum spin systems.
+By deriving universal scaling forms, we successfully captured the numerical data and extracted the critical exponent $\nu$ via data collapse near (2+1)D QCPs.
+Importantly, by definition, the DRTE provides a clear signature to distinguish weakly first-order phase transitions and deconfined quantum critical points even in small sizes, which is highly challenging to detect using conventional local observables.
+Notably, RTE and DRTE are universally applicable, as their definitions do not depend on specific models or order parameters.
+They also have the potential to characterizing phase transitions that are difficult to probe with conventional local observables, such as topological transitions, and may reveal new universal critical behaviors.
+Our work establishes a general and effective framework for studying quantum phase transitions across a wide range of classes, including continuous, first-order, and particularly elusive weakly first-order transitions.
 Acknowledgements.
-We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang.
+We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang and Wenan Guo.
 Z.W.
@@
 In this work, we focus on RTE.
-From the above definitions, it is evident that the R’enyi entropy can be expressed as the ratio of partition functions at two different temperatures: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}$ (for simplicity, we consider $S^{(2)}$ in this work).
+From the above definitions, it is evident that the Rényi entropy can be expressed as the ratio of partition functions at two different temperatures: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}$ (for simplicity, we consider $S^{(2)}$ in this work).
 The partition functions can be calculated using the bipartite reweight-annealing algorithm [29, 33, 67, 68] in conjunction with stochastic series expansion (SSE) QMC simulations [36, 17, 37, 38, 40, 39].
@@
 But $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can not be obtained.
-The antidote comes from some known reference point $Z(n\beta,J)/Z^{n}(\beta,J)$.
-Considering a zero-temperature system and assuming the degeneracy of the ground state is $de$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln de$+$c((J-J_{c})L^{(1/\nu)})$, where $c$ represents the critical corrections, $J_{c}$ represents the critical point, and $\nu$ is the critical exponent of the correlation length.
-Therefore, we can use some states with known degeneracy as reference points, such as product states and discrete symmetry-breaking states.
+The antidote comes from some known reference point $Z(n\beta,J)/Z^{n}(\beta,J)$: [Assuming that we have calculated the ratios $Z(n\beta,J^{\prime})/Z(n\beta,J)$ and $Z(\beta,J^{\prime})/Z(\beta,J)$ using the method described above, the ratio $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can be obtained if $Z(n\beta,J)/Z^{n}(\beta,J)$ is known.
+For a given zero-temperature phase, and assuming the ground state degeneracy is $d_{e}$, one can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z^{n}(\beta)}=\ln de$.
+Therefore, we can use states with known degeneracies — such as dimer product states (with degeneracy 1 in this work) and discrete symmetry-breaking states (with degeneracy 2 in this work) — as reference points at a given parameter $J$.
+In this way, $Z(n\beta,J)/Z^{n}(\beta,J)$ can be determined by solving: $\frac{1}{1-n}\ln\frac{Z(n\beta,J)}{Z^{n}(\beta,J)}=\ln de$.]
+One might be concerned about how to deal with a Hamiltonian without a state with known degeneracy.
+In this way, we can introduce a temperature axis starting from infinite temperature and annealing down to the target temperature.
+The degeneracy at infinite temperature is known as $2^{N}$ (consider an N-spin system with spin-1/2).
 The RTE derivative of $J$ is
@@
 Similarly, we can obtain
-$\begin{split}\frac{dZ(n\beta))/dJ}{Z(n\beta))}&=-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta))}\end{split}$ (S6)
+$\begin{split}\frac{dZ(n\beta))/dJ}{Z(n\beta))}&=-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}\end{split}$ (S6)
 Therefore, the RTE derivative can be rewritten as
-$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta))}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}\bigg{]}$ (S7)
+$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}\bigg{]}$ (S7)
 The equations above is general and doesn’t depend on detailed quantum Monte Carlo (QMC) methods.
@@
 where $n_{J_{1}}$ means the number of $J_{1}$ operators including both diagonal and off-diagonal ones.
-It’s worth noting that there is no “$2$” anymore in the $\langle...\rangle_{Z(n\beta)}$ term, because $n_{J_{1}}$ here contains two replicas’ operators which has already been doubled actually.
+It’s worth noting that there is no “22” anymore in the $\langle...\rangle_{Z(n\beta)}$ term, because $n_{J_{1}}$ here contains two replicas’ operators which has already been doubled actually.
+Note that the RTE is computed using the reweight-annealing QMC method mentioned above, while the DRTE is obtained directly by separately evaluating the two expectation values involved in Eq.
+(S9) or Eq.
+(S10).
+This implies that the computational cost of evaluating DRTE is roughly twice that of energy calculations.
+Typically, we perform on the order of one million Monte Carlo steps when computing the energy, and we have adopted the same number of steps in the present study.
+Figure S1: $J$-$Q$ models, in which the antiferromagnetic Heisenberg interaction ($J$-term) acts on all nearest-neighbor bonds.
+(a) The checkerboard $J$-$Q$ model on a square lattice with periodic boundary condition in both directions.
+The four-spin interaction ($Q$-term) acts on nearest-neighbor bond pairs that form a plaquette.
+(b) $J-Q_{2}$ model: the four-spin $Q$ interaction covers the entire lattice.
+(c) $J-Q_{3}$ model: the six-spin $Q$ interaction covers the entire lattice.
+Appendix B JQJ-Q models
+In this work, we studied three types of $J$-$Q$ models, namely checkerboard $J$-$Q$, $J$-$Q_{2}$, and $J$-$Q_{3}$ models.
+The Hamiltonian of the checkerboard $J$-$Q$ model on a square lattice is given by [49]
+$H=-J\sum_{\langle ij\rangle}P_{ij}-Q\sum_{[ijkl]}P_{ij}P_{kl},$ (S11)
+[…]
+$H=-J\sum_{\langle ij\rangle}P_{ij}-Q\sum_{[ijklmn]}P_{ij}P_{kl}P_{mn},$ (S12)
+, as illustrated in Fig. S1(c). The DRTE results for three models are shown in Figs. S2, along with the derivatives of the logarithmic partition function in panels (b,c), (e,f) and (h,i).
+Figure S2: We set $Q=1$ and use $J$ as the tuning parameter.
+(a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$.
+An inset shows the data on the right side of the phase transition point.
+(b) and (c) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model.
+(d–f) Data for the $J-Q_{3}$ model.
+(g–i) Data for the $J-Q_{2}$ model.

### v3

@@
-Extracting the singularity of a logarithmic partition function to investigate quantum phase transitions
+Complete finite-size scaling theory of Rényi thermal entropy for second, first and weak first order quantum phase transitions
 Abstract
-In principle, the logarithm of a partition function and its derivatives can be used to rigorously determine the nature of phase transitions.
-However, this approach is often ineffective in numerical calculations in practice, since resolving the singularity of a logarithmic partition function requires extremely high-precision data—especially when distinguishing weakly first-order phase transitions and deconfined quantum critical points.
-Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the analytic terms.
-This is achieved by computing the Rényi thermal entropy (RTE) and its derivative (DRTE).
-We have derived the scaling relations for RTE and DRTE, which successfully determine the phase transition points and critical exponents through data collapse in various quantum many-body systems.
-Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in JQJ-Q models, which were initially thought to be continuous deconfined quantum phase transitions.
-The discontinuity of these transitions is challenging to detect using traditional local physical quantities.
-This work provides a new and powerful paradigm for studying quantum many-body phase transitions.
-Introduction.— Quantum phase transitions, associated spontaneous symmetry breaking, and universality classes are central concepts in condensed matter and statistical physics [1, 2, 3].
-The nature of a phase transition is determined by the singularity in the free energy or its derivatives: a first-order transition is marked by a discontinuity in the first derivative, while a continuous transition involves singular higher-order derivatives.
-In principle, an exact calculation of the partition function would allow one to identify the transition based on these singularities.
-However, in practice, detecting such features in numerical simulations is challenging, as it requires extremely high-precision data.
-This issue is particularly severe when distinguishing weakly first-order transitions from deconfined quantum critical points [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
-A key difficulty arises because the singular contribution is often overwhelmed by a much larger, non-singular term that scales with the system volume.
-Extracting the singular part therefore remains a major challenge in computational physics.
-For a quantum phase transition (inverse temperature $\beta\rightarrow\infty$; in the realistic simulations, the finite-size system gap $\Delta_{L}\sim\frac{1}{L^{z}}$, $\beta\sim L^{z}$ is usually used for probing the ground state properties, in which $L$ is the system size and $z$ is the dynamical critical exponent [17] ), the singularity of the free energy can be equivalently studied through the logarithm of the partition function.
+Establishing the nature of a quantum phase transition in finite-size simulations – whether continuous, first-order, or weak first-order – is a fundamental challenge in quantum many-body computation.
+Especially, the weak first-order phase transition is affected by a super large correlation length and always displays as a continuous critical point in simulated finite-sizes.
+The core difficulty lies in the fact that there is no effective finite-size theory to distinguish these phase transitions in the realistic simulations limited by the computational resource.
+In this work, we have fixed this problem by introducing a unified finite-size framework based on the Rényi thermal entropy (RTE) and its derivative (DRTE) to detect and characterize quantum phase transitions.
+We derive complete scaling theories for the RTE and DRTE at second-order, first-order, and weak first-order transitions, showing that the DRTE naturally isolates the singular part of the free energy and strengthens the characteristics of various phase transitions in finite sizes.
+Using quantum Monte Carlo simulations, we demonstrate accurate data collapse and extraction of critical exponents at (2+1)-dimensional O($N$) critical points.
+More importantly, the DRTE provides a smoking-gun signature of weak first-order transitions through a clear double-peak structure and a crossing at zero, which we unambiguously observe in debated deconfined quantum criticality candidates such as the JJQQ models.
+Our approach offers a general, unbiased, and numerically efficient tool for probing the universal properties of quantum phase transitions, resolving long-standing ambiguities between continuous and weak first-order scenarios.
+I Introduction
+Quantum phase transitions, associated spontaneous symmetry breaking, and critical behaviors are central concepts in condensed matter and statistical physics [1, 2, 3].
+The nature of a phase transition is determined by singularities in the free energy or its derivatives, as rigorously defined in standard textbooks: a first-order transition is characterized by a discontinuity in the first derivative of the free energy, while a continuous transition involves singular higher-order derivatives.
+In principle, an exact analytical calculation of the partition function would allow unambiguous identification of the transition type through these singularities.
+However, only a handful of systems admit exact solutions, and the vast majority must be studied numerically.
+In practice, detecting such singular features in numerical simulations is highly challenging, as it requires extremely high-precision data to resolve the behavior of derivatives—particularly in strongly correlated many-body systems.
+A long-standing and fundamental challenge lies in identifying genuine deconfined quantum critical points (DQCP) beyond the Landau-Ginzburg-Wilson paradigm.
+This framework describes continuous transitions between incompatible symmetry-breaking phases, such as the antiferromagnetic (AF) Néel order—where spin symmetry is broken—and the valence-bond solid (VBS)—where lattice-rotation symmetry is broken [4, 5, 6].
+Such continuous transitions exhibit unconventional critical behavior and may host emergent symmetries and fractionalized excitations [7, 8, 9, 10, 11, 12, 13, 14].
+However, in concrete model realizations, it remains controversial whether candidate DQCPs represent true continuous transitions or are instead weak first-order.
+Notable examples include nonlinear sigma models with topological terms [15, 16, 17, 18], the JQJ-Q models [5, 19, 20, 21, 22, 23], and other bosonic or fermionic systems [24, 25, 26, 27].
+The core difficulty stems from the fact that correlation lengths at these candidate DQCPs are often much longer than the system sizes accessible in simulations—even if the transition is weak first-order rather than genuinely critical.
+Consequently, the observed critical-like behavior closely mimics that of a continuous transition, making definitive discrimination extremely difficult.
+This phenomenon is understood within a physical picture in which the weak first-order transition occurs in close proximity to a critical point—though some researchers argue that the underlying criticality is described by a complex conformal field theory (CFT)[28, 29, 18, 30], while others propose it arises from a multicritical point[31, 32, 33].
+As a result, in finite-size simulations, weak first-order transitions near a critical fixed point are nearly indistinguishable from continuous ones.
+Given current computational limits, reliably distinguishing between these scenarios across accessible system sizes remains a major challenge.
+Crucially, there is currently no unified and effective finite-size scaling theory capable of robustly differentiating second-order and weak first-order transitions.
+As a consequence, most numerical studies lack a definitive “smoking-gun” signature for identifying the true nature of the transition.
+In this work, we address this gap by developing a complete scaling theory based on the Rényi thermal entropy (RTE) and its derivative (DRTE)—physical observables that naturally cancels the leading analytic contributions to the free energy, thereby isolating the singular part that governs critical behavior.
+Moreover, the DRTE exhibits distinct and universal signatures in both first-order and weak first-order transitions, enabling clear discrimination between different types of quantum phase transitions.
+In continuous and weak first-order cases, it allows for the accurate extraction of critical exponents through data collapse, where the extracted exponents in the latter reflect those of the nearest critical fixed point.
+Thus, our framework enables the nature of quantum phase transitions to be unambiguously identified through finite-size scaling of the RTE and DRTE, providing a general, unbiased, and numerically efficient tool for resolving long-standing debates in quantum criticality.
+II Second order phase transition
+II.1 Finite-size scaling form
+Let us start from the second order phase transition.
+Since we discuss quantum phase transition, $\beta\sim L^{z}$ is required ($\beta=1/T$ is the inverse temperature, $L$ is the system size and $z$ is the dynamical critical exponent).
+The free energy $F=-\frac{1}{\beta}lnZ$ of a system can be replaced by a logarithm of the partition function $lnZ$ because $\beta$ is irrelevant here.
+As we know, due to the correlation length diverging at a second order phase transition point, the free energy $F$ or the logarithmic partition function $lnZ$ must have a singular term.
 According to the finite-size scaling hypothesis [34, 35], a singular observable $Q$ (not necessarily divergent) near the critical point $g_{c}$ scales with system size as:
 $Q(\Delta g,L)=L^{\sigma}\tilde{S}(\Delta gL^{1/\nu})$ (1)
-where $\Delta g=g-g_{c}$, $\nu$ is is the correlation length critical exponent, $\sigma$ is a scaling power and $\tilde{S}$ is a universal scaling function.
+where $\Delta g=g-g_{c}$, $g$ is the tunable parameter of a quantum phase transition and $g_{c}$ is the parameter at the critical point.
+$\nu$ is the correlation length critical exponent, $\sigma$ is a scaling power and $\tilde{S}$ is a universal scaling function.
 Since the singular part of the free energy is scale-invariant ($\sigma=0$[36, 34], the logarithm of the partition function can be expressed as:
 $\ln Z(g,\beta,L)=a(g)\beta L^{d}+\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}(\Delta gL^{1/\nu})$ (2)
-where $Z$ is a partition function, $d$ is the spatial dimension, $L$ is the system length, inverse temperature $\beta\sim L^{z}$ and $a(g)\beta L^{d}$ represents the leading term obeying a volume law, and $\mathcal{A}_{sub}(g,L,\beta)=\beta(a_{1}(g)L^{d-1}+a_{2}(g)L^{d-2}+a_{3}(g)L^{d-3}+\cdots)$ accounts for sub-leading analytic contributions ($a_{i}$ can be zero or non-zero).
-The dominant volume-law term often obscures the singular contribution $\tilde{S}$, which explains the difficulty in detecting phase transitions numerically through such singularities.
+where $Z$ is a partition function, $d$ is the spatial dimension, $L$ is the system length, inverse temperature $\beta\sim L^{z}$ and $a(g)\beta L^{d}$ represents the leading term obeying a volume law, and $\mathcal{A}_{sub}(g,L,\beta)=\beta(a_{1}(g)L^{d-1}+a_{2}(g)L^{d-2}+a_{3}(g)L^{d-3}+\cdots)$ accounts for sub-leading analytic contributions ($a_{i}$ can be zero).
+The dominant analytic terms often obscure the singular contribution $\tilde{S}$, which explains the difficulty in detecting phase transitions numerically through such singularities.
 To isolate the singular term $\tilde{S}$, one can construct a quantity in which the dominant analytic contributions cancel out.
@@
 Specifically, we argue that Eq.
-(2) can be rewritten as $\ln Z(2\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}$ and $2\ln Z(\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+2\tilde{S}$.
-Coincidentally, this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [22, 23, 24, 25, 26, 27, 28]:
+(2) can be rewritten as
+$\ln Z(2\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}$ (3)
+and
+$2\ln Z(\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+2\tilde{S}$, (4)
+so that their difference $\ln Z(2\beta)-2\ln Z(\beta)]$ scales as $\sim\tilde{S}$.
+This is indeed physically reasonable, because for a quantum phase transition, temperature is an irrelevant field.
+As long as it is proportional to the finite-size energy gap ($\beta\sim L^{z}\rightarrow\infty$), different multiples should only affect the analytic terms.
+We originally intended to assign a name to the difference between the two physical quantities.
+Coincidentally, this difference turns out to be exactly the second-order Rényi thermal entropy—also known as the purity—a fundamental quantity in information theory [39, 40, 41, 42, 43, 44, 45]:
 $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$ (5)
@@
 For clarity and simplicity, we focus on the second-order RTE ($n=2$) throughout this work.
-In this work, using our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [29, 30, 31, 32, 33, 34, 35], we validate the conjecture that the Rényi thermal entropy (RTE) is a dimensionless quantity capturing only the singular contribution.
-Moreover, its derivative (DRTE) exhibits a more pronounced singularity near the quantum critical point (QCP).
-We propose universal scaling forms for both RTE and DRTE, which are confirmed through data collapse analyses at (2+1)-dimensional (D) QCPs in the O($N$) universality classes, with $1\leq N\leq 3$.
-Notably, because the leading non-singular term in $\ln Z$ is canceled in DRTE, it provides a powerful diagnostic for weakly first-order phase transitions, even in relatively small systems (up to $L=56$), which are significant challenges to resolve using conventional numerical approaches.
-Models and method.— We investigate quantum spin models with continuous QCPs in the (2+1)D O($N$) universality classes, where $1\leq N\leq 3$.
-We also examine first-order transitions, with a focus on weakly first-order cases, including some debated deconfined quantum critical points.
-The model Hamiltonians are described in detail in the following sections.
-The second-order RTE can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$.
-We compute the partition functions using the bipartite reweight-annealing algorithm [29, 30] in combination with stochastic series expansion (SSE) QMC simulations [36, 17, 37, 38, 39, 40, 41].
-Following a method analogous to that used for the derivative of Rényi entanglement entropy and negativity [29, 42], the derivative of the RTE with respect to a general parameter $J$ is given by:
-$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
-Here, the two expectation values are evaluated separately using ensembles corresponding to $Z(n\beta)$ and $Z(\beta)$, respectively.
-While the RTE is calculated using the reweight-annealing QMC method, the derivative (DRTE) is directly obtained from the two expectation values in Eq.
-(4); see Supplementary Materials (SM) for details.
-Considering the dynamical critical exponents $z=1$ of the models studied in this work, the $\beta=L$ is set to approach the quantum critical regime.
-Finite-size scaling form.—For a zero-temperature system with ground-state degeneracy $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$$\tilde{S}(\Delta gL^{1/\nu})$, where $\tilde{S}$ represents the critical corrections.
-As shown in Fig.1, $S^{(2)}$ effectively probes both the degeneracy of discrete symmetry-breaking phases and the tower of excited states (TOS) associated with continuous symmetry-breaking.
-Near criticality, the logarithmic partition function consists of a dominant non-singular term $a(J)\beta L^{d}$ and a dimensionless singular term $\tilde{S}(\Delta gL^{1/\nu})$.
-The leading non-singular term obscures the critical behavior, making it difficult to extract the singular part directly from the partition function.
-More interestingly, as shown in Fig.
-1, the curves of $S^{(2)}$ for different system sizes intersect at the critical point, indicates that the RTE is independent of system size, providing strong numerical evidence that the leading non-singular contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel exactly in the definition of RTE.
-This cancellation isolates the dimensionless singular term.
-Moreover, the DRTE displays a more pronounced divergence near the QCPs, further highlighting its sensitivity to critical fluctuations.
-According to the finite-size scaling hypothesis  [18, 17], any singular quantity near a critical point follows a scaling form as given in Eq.(1).
-In particular, the dimensionless nature of the RTE ($\sigma=0$), implies the scaling form:
+In particular, according to Eq.(1), the dimensionless nature of the RTE ($\sigma=0$), implies the scaling form:
 $S^{(n)}(\Delta g,L)=\tilde{S}(\Delta gL^{1/\nu})$ (6)
-From Eq. (5), we can readily derive the scaling behavior of the DRTE:
-$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
+From Eq. (6), we can readily derive the scaling behavior of the derivative of RTE (DRTE):
+$\frac{\partial S^{(n)}(\Delta g,L)}{\partial\Delta g}=\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (7)
 where $\tilde{S}^{\prime}(x)$ is $\partial\tilde{S}(x)/\partial x$ and $x$ is a general parameter.
+If our above analysis is correct, the RTE should have a crossing at a critical point since it is dimensionless like Binder ratio.
+And the DRTE can extract the critical exponent $\nu$ by data collapse.
+In the following section, by expanding the scaling function $\tilde{S}(\Delta gL^{1/\nu})$ (or $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$) into polynomials: $\tilde{S}(x)=\sum_{k=0}^{k_{\mathrm{max}}}\frac{1}{k!}\tilde{S}_{0}^{(k+1)}(0)x^{k}$.
+we will numerically prove our scaling formula near the widely accepted (2+1)D O(1) ($Z_{2}$), O(2), and O(3) quantum critical points (QCPs) through perfect data collapse in Fig.1.
+The critical point $g_{c}$, the critical exponent $\nu$, and the coefficients $\tilde{S}_{0}^{(k+1)}(0)$ are the fitting parameters and $k_{\mathrm{max}}=6$ in the following numerical analysis.
 Figure 1: (a,c) Second-order RTE and its derivative DRTE, and (b,d) corresponding data collapse analysis near the QCP of the bilayer Ising–Heisenberg model [(2+1)D Ising universality].
@@
 To achieve good data collapse, we include a subleading correction: $Q(g,L)=L^{k}[\tilde{S}((g-g_{c})L^{1/\nu})+bL^{-\omega}]$.
-$\omega$ is close to 1, and to reduce the numerical instability caused by increasing the number of fitting parameters, we set $\omega=1$.
+$\omega=0.830$, $0.789$, and $0.782$ correspond to the (2+1)D Ising, O(2), and O(3) critical points, respectively [46, 47, 48, 49].
 For $Q=S^{(2)}$, $k=0$.
 $k=1/\nu$ when $Q$ is DRTE.
-In the following section, by expanding the scaling function $\tilde{S}(\Delta gL^{1/\nu})$ (or $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)D O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
-Ising transition.— We study the bilayer Ising-Heisenberg model with a (2+1)D Ising QCP [43]. The Hamiltonian is given by:
+II.2 Numerical methods
+The RTE can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$.
+We compute the partition functions using the bipartite reweight-annealing algorithm [50, 51, 52] in combination with stochastic series expansion (SSE) QMC simulations [53, 35, 54, 55, 56, 57].
+Following a method analogous to that used for the derivative of Rényi entanglement entropy and negativity [50, 58], the derivative of the RTE with respect to a general parameter $g$ of the Hamiltonian $H$ is given by:
+$\frac{dS^{(n)}}{dg}=\frac{1}{1-n}\bigg[-n\beta\bigg\langle\frac{dH}{dg}\bigg\rangle_{Z(n\beta)}+n\beta\bigg\langle\frac{dH}{dg}\bigg\rangle_{Z(\beta)}\bigg]$ (8)
+Here, the two expectation values are evaluated separately using ensembles corresponding to $Z(n\beta)$ and $Z(\beta)$, respectively.
+While the RTE is calculated using the reweight-annealing QMC method [59, 60, 61], the DRTE is directly obtained from the two expectation values in Eq.
+(8); see Appendix A for details.
+Considering the dynamical critical exponents $z=1$ of the models studied in this work, the $\beta=L$ is set to approach the quantum critical regime.
+II.3 Ising transition
+We study the bilayer Ising-Heisenberg model with a (2+1)D Ising QCP [62]. The Hamiltonian is given by:
 $H=-J_{z}\sum_{\alpha=1,2}\sum_{\langle ij\rangle}S_{i,\alpha}^{z}S_{j,\alpha}^{z}+J\sum_{i}{\mathbf{S}}_{i,1}\cdot{\mathbf{S}}_{i,2},$ (9)
@@
 (6) and (7) to the RTE and DRTE data, respectively.
-The fits yield $J_{z,c}=0.9999(1)$, $\nu=0.62(1)$ from RTE, and $J_{z,c}=1.000(3)$, $\nu=0.63(4)$ from DRTE.
-These values are agreement with the known QCP $J_{z,c}=1$  [43] and the (2+1)D Ising universality class value $\nu=0.63012$  [44, 45, 43].
-The resulting universal scaling functions $\tilde{S}(\Delta gL^{1/\nu})$ and $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$, obtained from data collapse, are displayed in Figs.
+The fits yield $J_{z,c}=0.9999(1)$, $\nu=0.629(7)$ from RTE, and $J_{z,c}=1.000(3)$, $\nu=0.63(4)$ from DRTE.
+These values are in agreement with the known QCP $J_{z,c}=1$  [62] and the (2+1)D Ising universality class value $\nu=0.63012$  [63, 64, 62].
+The resulting universal scaling functions $\tilde{S}(\Delta gL^{1/\nu})$ and $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$ with $\Delta g=J_{z}-J_{z,c}$, obtained from data collapse, are displayed in Figs.
 1(b) and (d), respectively.
-Continuous symmetry breaking transitions.— We then investigate QCPs of the dimerized AF Heisenberg model [46, 47, 48]. The Hamiltonian reads:
+II.4 Continuous symmetry phase transitions
+We then investigate QCPs of the dimerized AF Heisenberg model [65, 66, 67]. The Hamiltonian reads:
 $H=J_{1}\sum_{\langle ij\rangle}D_{ij}+J_{2}\sum_{\langle ij\rangle^{\prime}}D_{ij},$ (10)
@@
 The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $0<\Delta<1$/ $\Delta>1$ / $\Delta=1$  [65, 66, 67].
+For a zero-temperature system with ground-state degeneracy $d_{e}$, the RTE actually reflects the logarithmic degeneracy: $S^{(2)}=\ln d_{e}$.
+As shown in Fig.1, $S^{(2)}$ effectively probes both the degeneracy of discrete symmetry-breaking phases and the Anderson tower of states associated with continuous symmetry-breaking [68, 69, 12].
 We evaluate the RTE and its derivative near the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality), with results shown in Figs.
@@
 Similar to the Ising case, the RTE curves for different system sizes intersect at the QCPs, and their derivatives exhibit clear peaks at these points.
-Figure 2: We set $Q=1$ and use $J$ as the tuning parameter.
-(a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$.
-The inset shows data on the right side of the transition point.
-(b) and (c) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model.
-(d–f) Data for the $J-Q_{3}$ model.
-(g–i) Data for the $J-Q_{2}$ model.
-The raw values of the partition function derivative vary significantly across system sizes (see SM), making direct comparison difficult.
-To resolve this, we normalize the data by $L^{3}$, corresponding to $\beta L^{2}=L^{3}$.
 We apply the scaling forms in Eqs.
 (6) and (7) to data near the (2+1)D O(2) and O(3) QCPs.
-For the O(2) QCP, fitting the RTE yields $J_{1,c}=0.997(1)$, $\nu=0.67(2)$, while the DRTE gives $J_{1,c}=0.999(1)$, $\nu=0.66(2)$.
+For the O(2) QCP, fitting the RTE yields $J_{1,c}=0.997(1)$, $\nu=0.67(2)$, while the DRTE gives $J_{1,c}=0.9994(8)$, $\nu=0.66(1)$.
 These values are consistent with the known critical parameters $J_{1,c}=1$ and $\nu=0.6703$ [63, 67].
-For the O(3) QCP, fitting the RTE data yields $J_{1,c}=0.523(1)$, $\nu=0.70(1)$, while the DRTE gives $J_{1,c}=0.525(1)$, $\nu=0.71(2)$.
+For the O(3) QCP, fitting the RTE data yields $J_{1,c}=0.523(1)$, $\nu=0.703(7)$, while the DRTE gives $J_{1,c}=0.525(1)$, $\nu=0.708(8)$.
 These values are consistent with the known critical parameters $J_{1,c}=0.52337$ and $\nu=0.7073$ [63, 65, 66].
@@
 1(f,h) and (j,m), respectively.
-First-order and weakly first-order transitions.— Having established the universal scaling behavior at (2+1)D QCPs, we now turn to first-order phase transitions, with a focus on weakly first-order transitions, which presents significant challenges for traditional local physical quantities.
-We first study the 2D checkerboard $J$-$Q$ model ( see SM for model details), which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [49].
-The DRTE around this transition is plotted in Fig.
-2 (a).
-A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side as $L\to\infty$.
-This behavior arises because DRTE isolates the singular contribution to the logarithmic partition function (as mentioned above).
-For comparison, we also calculate the derivative of logarithmic partition functions, as shown in Figs.
-2 (b) and (c).
-These curves appear much smoother across the transition compared with DRTE, due to the dominance of a non-singular, volume-law term.
-The phase transition of this model is obviously first-order according to the previous works [49, 29], thus we see that the curves in (b) and (c) have very small kinks around the phase transition point, especially for larger sizes.
-The phase transition between AF Néel order and valence-bond solid (VBS) phase in 2D quantum magnets can appear (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [4, 5].
+In summary, all the data show that the RTE does have a crossing at a critical point and the DRTE does extract the critical exponent $\nu$ by data collapse, which is highly consistent with our theoretical analysis.
+III First-order phase transition
+III.1 Finite-size scaling form
+At a first-order phase transition point $g_{c}$, the free energy densities of the two phases, $f_{1}(g)$ and $f_{2}(g)$, are equal, but their slopes differ (i.e., $\Delta f^{\prime}(g_{c})\neq 0$).
+Near the first-order phase transition point, for a finite system of linear size $L$ and spatial dimension $d$, at sufficiently low temperatures, the partition function of the system can be approximated as a simple sum of contributions from the two phases [70]: $Z(g,\beta,L)\simeq e^{-\beta L^{d}f_{1}(g)}+e^{-\beta L^{d}f_{2}(g)}.$ Here, contributions from the interface between the two phases and corrections due to fluctuations are neglected; this approximation holds when $L$ is large and the temperature is low.
+Each term corresponds to the free energy of a pure phase, $F_{i}=L^{d}f_{i}(g)$.
+The logarithm of the partition function can be expressed as $\ln Z(g,\beta,L)=-\beta L^{d}f_{1}(g)+\ln\!\left[1+e^{-\beta L^{d}\Delta f(g)}\right]$, where we define the free energy density difference as $\Delta f(g)=f_{2}(g)-f_{1}(g)$.
+From this expression, we can further derive the second Rényi entropy as
+$S^{(2)}=2\ln\!\left(1+e^{-\beta L^{d}\Delta f}\right)-\ln\!\left(1+e^{-2\beta L^{d}\Delta f}\right).$ (11)
+Thus, the DRTE takes the compact expression
+$\frac{\partial S^{(2)}}{\partial g}=2\beta L^{d}\,\Delta f^{\prime}(g)\bigl[w(2\beta)-w(\beta)\bigr],$ (12)
+where $w(\beta)=\frac{1}{1+e^{-\beta L^{d}\Delta f(g)}}$ is the logistic function, which here represents the statistical weight associated with one of the two phases.
+Figure 2: (a) Derivative of the RTE, and (b) corresponding data collapse analysis near the phase transition point of the 2D anisotropic Heisenberg model with anisotropy parameter $\Delta$ ( see Eq.
+(10)).
+(c) and (d) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition.
+Here, $J_{1}=J_{2}=1$, and the anisotropy $\Delta$ is taken as the tuning parameter.
+(e–h) Results for the checkerboard $J$-$Q$ model, using $Q=1$ and tuning $J$.
+The Eq. (12) immediately explains 3 characteristic features of the DRTE at a first-order transition:
+(1) Since $\Delta f^{\prime}(g_{c})\neq 0$, the difference $w(2\beta)-w(\beta)$ changes sign across $g_{c}$ (near $g_{c}$, $\Delta f(g)\approx\Delta f^{\prime}(g_{c})(g-g_{c})$, crossing zero linearly), resulting in a robust double-peak structure with one positive and one negative peak located on opposite sides of the transition.
+This discontinuous structure reflects the rapid transfer of thermodynamic weight between the two competing phases and is a direct consequence of phase coexistence.
+Indeed, this is fully consistent with the definition of a first-order phase transition.
+The RTE is strictly defined as the difference in the logarithm of the partition function at different inverse temperatures: $S^{(n)}=n\ln Z(\beta)-\ln Z(n\beta)$, where different values of $n$ correspond to different linear rates of approaching zero temperature.
+For quantum phase transitions, the inverse temperature $\beta$ is a non-relevant field, and thus the difference in the logarithm of the partition function corresponds to the difference in free energy.
+Mathematically, the sum or difference of two continuous functions remains continuous.
+In other words, if the resulting function is discontinuous, then at least one of the original functions must be discontinuous.
+In this case, the DRTE is such a resultant function, and its discontinuity directly reflects a discontinuity in the first derivative of the free energy itself.
+Therefore, the discontinuity in the DRTE directly characterizes a first-order phase transition.
+(2) Due to $\Delta f(g_{c})=f_{2}(g_{c})-f_{1}(g_{c})=0$, it follows that $w(2\beta)-w(\beta)=0$, and therefore the DRTE equals zero at the transition point.
+This represents another fundamental distinction from continuous phase transitions.
+In continuous transitions, the divergence of the correlation length leads to a critical singularity, resulting in a divergent behavior of the DRTE that scales with system size as $L^{1/\nu}$ (as shown in Eq.
+(7)).
+(3) Because $|w(2\beta)-w(\beta)|$ is an $O(1)$ constant, and $\Delta f^{\prime}(g)$ is also of constant order, it follows that the peak amplitude of the DRTE scales as
+$\left|\partial S^{(2)}/\partial g\right|\sim\beta L^{d}\sim L^{d+z}.$ (13)
+For quantum phase transitions in $(2+1)$ dimensions with dynamical exponent $z=1$ (so that $\beta\sim L$), gives $\left|\partial S^{(2)}/\partial g\right|\sim L^{3}.$ For a first-order phase transition, the inverse of the correlation length exponent is given by $1/\nu=d+z$ effectively (here, $1/\nu=3$).
+It implies that DRTE satisfies the following finite-size scaling ansatz:
+$\left|\partial S^{(2)}/\partial g\right|=L^{d+z}f(\Delta gL^{d+z}).$ (14)
+It is worth noting that the DRTE characterizes the difference in free energy, rather than the difference in free energy density.
+This divergence does not arise from a critical singularity, but rather stems from the intrinsic volume-law scaling of the system and the fact that the free energies of the coexisting phases are no longer equal once the system moves away from the transition point.
+III.2 Numerical demonstration
+Our numerical simulations confirm our theoretical derivation.
+Specifically, we study the 2D anisotropic Heisenberg model and the 2D checkerboard $J$-$Q$ (CBJQ) model [71].
+Let us begin by examining the 2D anisotropic Heisenberg model defined in Eq.
+(10)), where the antiferromagnetic couplings are uniform, $J_{1}=J_{2}=1$, and the anisotropy parameter $\Delta$ serves as the tuning variable.
+At $\Delta=1$, the system exhibits full spin-rotation symmetry, corresponding to O(3), which is spontaneously broken by the emergence of AF order in the ground state.
+When $\Delta$ deviates from unity, the symmetry is explicitly lowered to in-plane spin rotation symmetry and spin-flip symmetry.
+For $0<\Delta<1$, the ground state develops an easy-plane AF order, spontaneously breaking the continuous $\mathrm{SO}(2)$ symmetry.
+In contrast, for $\Delta>1$, the system favors an easy-axis AF order, leading to the breaking of the discrete $\mathbb{Z}_{2}$ spin-inversion symmetry.
+Precisely at $\Delta=1$, both types of order coexist, and each vanishes discontinuously across the transition.
+This abrupt change in the order parameters signals a first-order quantum phase transition, accompanied by the spontaneous breaking of an enhanced O(3) symmetry [71].
+The DRTE across the transition at $\Delta_{c}=1$ is shown in Fig.
+2(a).
+A clear discontinuity is observed near the phase transition point, with the DRTE diverging toward opposite signs on either side of the transition in the thermodynamic limit ($L\to\infty$).
+For comparison, we also compute the derivative of the logarithm of the partition function, as displayed in Figs.
+2(c) and (d).
+These curves exhibit a significantly smoother behavior across the transition compared to the DRTE.
+Moreover, the DRTE curves for different system sizes are seen to intersect precisely at the critical point $\Delta_{c}=1$, and the value at the crossing point is zero—consistent with the expected vanishing of the DRTE at a first-order transition.
+Finally, as shown in the data collapse in Fig.
+2(b), the DRTE obeys the scaling form given by Eq.
+(14).
+These results provide strong numerical confirmation of our theoretical predictions.
+The occurrence of a symmetry-enhanced first-order transition is not limited to systems with exact symmetries.
+An emergent enhanced symmetry can arise at long wavelengths provided that all explicit symmetry-breaking perturbations are irrelevant under the renormalization group (RG) and thus decay along the RG flow.
+In this scenario, we study the 2D checkerboard $J$-$Q$ model (see Appendix B for model details), which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$, where the system transitions from an O(3)-symmetry-breaking antiferromagnetic order to a $\mathbb{Z}_{2}$-symmetry-breaking plaquette-singlet solid order [71].
+The DRTE across this transition is shown in Fig.
+2(e).
+A clear discontinuity is observed at the critical point, with the DRTE diverging to opposite signs on either side of the transition in the thermodynamic limit ($L\to\infty$).
+The RTE equals zero at the transition point.
+Furthermore, the DRTE data collapse successfully onto a single curve when rescaled according to Eq.
+(14) (Fig.
+2(f)), consistently confirming the validity of our theoretical predictions.
+Similar to the anisotropic Heisenberg model, the derivative of the logarithm of the partition function (shown in Figs.
+2(g) and (h)) exhibits a much smoother behavior across the transition compared to the DRTE.
+Combining theoretical analysis with numerical simulations, we demonstrate that the DRTE is not only effective for characterizing continuous phase transitions, but also serves as a powerful probe for first-order phase transitions.
+IV weak first-order transitions
+IV.1 Finite-size scaling form
+Having established the universal behaviors at second order QCPs and first order phase transitions, we now turn our attention to weak first order transitions, which pose significant challenges for conventional local physical observables.
+The fundamental challenge is that there is a very large correlation length at the phase transition point, thus all the behaviors in finite sizes are similar to a critical point.
+The weak first order phase transitions are generally believed to be governed by two nearby fixed points in finite sizes: one corresponding to the first order transition that the system flows into in the thermodynamic limit, and the other associated with a continuous phase transition in close proximity.
+For a continuous transition, the correlation length diverges near the critical point as $\xi\sim|J-J_{c}|^{-\nu}$.
+When these two fixed points are very close—i.e., when $|J-J_{c}|$ is small—the correlation length $\xi$ can become large even for a slight deviation from criticality.
+As a result, for finite system sizes with $L<<\xi$, the system lies within the critical region of the nearby continuous transition, even as it approaches the first-order transition point.
+Consequently, the system exhibits signatures of continuous critical behavior, despite the underlying transition being first-order.
+This crossover effect explains the observed scaling collapse using continuous-transition exponents in finite-size systems.
+Figure 3: We set $Q=1$ and use $J$ as the tuning parameter.
+(a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$.
+(b) Corresponding data collapse analysis of (a).
+(c) and (d) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the $J-Q_{3}$ model.
+(e–h) Data for the $J-Q_{2}$ model.
+The raw values of the partition function derivative vary significantly across system sizes (see Appendix B), making direct comparison difficult.
+To resolve this, we normalize the data by $L^{3}$, corresponding to $\beta L^{2}=L^{3}$.
+Let us go back to the scaling behavior of DRTE that captures characteristics of the standard first order transition in Eq.
+(12), where actually we have ignored the quantum fluctuations since it is a strong first order phase transition.
+However, in the case of a weak first-order transition, quantum fluctuations become essential and must be properly accounted for.
+In finite-size simulation, quantum fluctuations originating from a nearby critical point introduce an effective “singular” contribution—though not a true thermodynamic singularity—that mimics the behavior observed in second-order transitions (as shown in Eq.(7)).
+We therefore conjecture that the DRTE expression in Eq.(12) should be modified in the vicinity of such a critical point to include this effect:
+$\frac{\partial S^{(2)}}{\partial g}=(2L^{d+z}\,\Delta f^{\prime}(g)+\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu})\bigl[w(2\beta)-w(\beta)\bigr],$ (15)
+The Eq. (15) predicts some features of DRTE behaviors at a weak first order phase transition point:
+1) The double-peak structure in first order phase transition is still present, since the $w(2\beta)-w(\beta)$ term remains.
+2) The DRTE also equals zero at the transition point, since the fact of energy-crossing at first order phase transition is unchanged.
+3) The DRTE does not follow the finite-size scaling form for a strong first-order transition in Eq.
+(14) for system sizes smaller than the correlation length, but instead obeys the “singular” correction term scaling as $\sim L^{1/\nu}$, where $\nu$ reflects the critical exponent of the nearby critical point.
+For the part $(2L^{d+z}\,\Delta f^{\prime}(g)+\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu})$ in Eq.(15), we argue that the coefficient of the term $2L^{d+z}\Delta f^{\prime}(g)$, i.e., $2\Delta f^{\prime}(g)$ is extremely small around the weak first order phase transition point, thus it will suppress the $L^{d+z}$ scaling in finite sizes and emphasize the $\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}$ scaling term.
+While the simulating sizes become much larger than the correlation length, the $L^{d+z}$ scaling dominates and the “singular" term can be ignored due to $d+z>1/\nu$, the Eq.(15) goes back to Eq.(12).
+In the following, we will show the numerical data which strongly support our analysis above.
+IV.2 Numerical demonstration
+The phase transition between AF Néel order and VBS phase in 2D quantum magnets can appear (approximately) continuous, which is thought as the deconfined quantum criticality (DQC) [72, 73].
 $J-Q_{2}$ and $J-Q_{3}$ models are prototypical examples that realize such transitions in 2D quantum spin systems [5, 22].
@@
 However, the critical exponents observed in these models do not satisfy the conformal bootstrap bounds expected for SO(5) critical point [77].
-Recent works have increasingly suggested that this DQC scenario may instead correspond to a weakly first-order transition or as a nonunitary CFT  [7, 11, 57, 58, 12, 59, 60, 61, 8, 15, 62, 63, 64, 65, 16].
-In the following, we show that the DRTE provides a clear and direct signature that the transitions in the $J-Q_{2}$ and $J-Q_{3}$ models ( see SM for models’ details) are indeed first-order, even in rather small systems (up to $L=56$).
+Recent works have increasingly suggested that this DQC scenario may instead correspond to a weak first order transition  [78, 79, 80, 81, 82, 21, 18, 31, 83, 84, 85].
+Next, we will show that the DRTE behaviors provide a clear and direct signature that the transitions in the $J-Q_{2}$ and $J-Q_{3}$ models ( see Appendix B for models’ details) are indeed weak first order in finite sizes (up to $L=56$).
+The observed DRTE signals fully align with the three key features predicted in the previous subsection, offering compelling evidence against a genuine continuous transition.
 The transition points of the 2D square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [5, 22, 86, 31].
 The DRTE results for both models are shown in Figs.
-2 (d) and (g), along with the derivatives of the logarithmic partition function in panels (e,f) and (h,i).
-Due to the dominant background contribution, the derivative of the logarithmic partition function appears smooth across the transition, even for large system sizes, as seen in Figs.
-2(e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model.
-This shows that the derivative of the logarithmic partition function can not be directly reveal the weakly first-order nature.
-In contrast, the DRTE reveals clear discontinuities at the transition points, providing direct evidence that both transitions are (weakly) first-order.
-Notably, the $J-Q_{2}$ model exhibits stronger finite-size effects than $J-Q_{3}$, which has made its first-order nature more difficult to resolve in previous studies [63].
-Our results demonstrate that DRTE can unambiguously detect even weakly first-order transitions, in relatively small systems (up to $L=56$).
-Conclusion and discussions.— We proposed that the RTE isolates the singular contribution of the logarithmic partition function, making it highly sensitive near phase transitions.
-We systematically studied the RTE and its derivative DRTE around QCPs in 2D quantum spin systems.
-By deriving universal scaling forms, we successfully captured the numerical data and extracted the critical exponent $\nu$ via data collapse near (2+1)D QCPs.
-Importantly, by definition, the DRTE provides a clear signature to distinguish weakly first-order phase transitions and deconfined quantum critical points even in small sizes, which is highly challenging to detect using conventional local observables.
-Notably, RTE and DRTE are universally applicable, as their definitions do not depend on specific models or order parameters.
-They also have the potential to characterizing phase transitions that are difficult to probe with conventional local observables, such as topological transitions, and may reveal new universal critical behaviors.
-Our work establishes a general and effective framework for studying quantum phase transitions across a wide range of classes, including continuous, first-order, and particularly elusive weakly first-order transitions.
+3 (a) and (e), along with the derivatives of the logarithmic partition function in panels (c,d) and (g,h).
+The derivative of the logarithmic partition function appears smooth across the transition, even for large system sizes, as seen in Figs.
+3(c,d) for $J-Q_{3}$ and (g,h) for $J-Q_{2}$ model.
+This shows that the almost smooth derivative of the logarithmic partition function cannot directly reveal the weak first-order nature, as discussed above, since the logarithmic partition function contains large analytic contributions.
+In contrast, the DRTE reveals clear discontinuities with a double-peak character near the transition points (see Figs.
+3 (a) and (e)), as the feature 1) we predicted in the above subsection.
+Moreover, DRTE curves for different system sizes intersect at zero at the transition point (as the feature 2) we predicted above), indicating the absence of singular behavior—unlike in continuous phase transitions.
+These two features are fully consistent with our previous analysis of the DRTE at weak and standard first order transitions.
+Our results demonstrate that the DRTE can unambiguously detect the first-order nature of weak first-order phase transitions in relatively small systems (up to $L=56$), even in the $J$-$Q_{2}$ model with much stronger finite-size effects.
+In fact, its nature, whether continuous or first-order, remains under debate so far [31].
+[…]
+It is worth noting that, in current studies based on local order parameters, the observed behavior within accessible system sizes is predominantly governed by signatures of continuous phase transitions, making it difficult to unambiguously identify the underlying first-order nature.
+In contrast, our approach—using the DRTE—successfully captures features of both continuous critical fluctuations and first-order discontinuity even in relatively small systems (up to $L=56$).
+This provides a clear and compelling demonstration of the weak first-order character of the transition, marking, to our knowledge, the first such identification achieved at modest system sizes.
+V Conclusion and discussions
+We have established a unified and comprehensive scaling theory for the Rényi thermal entropy (RTE) and its derivative (DRTE) across all three fundamental types of phase transitions: continuous (second-order), strong first-order, and weak first-order.
+Our framework is predicated on the key insight that the DRTE naturally isolates the singular part of the free energy by canceling dominant analytic contributions.
+Crucially, we have derived distinct and diagnostic scaling behaviors for the DRTE in each regime:
+1. Continuous Phase Transitions: The DRTE scales as
+$\frac{\partial S^{(2)}}{\partial g}=L^{1/\nu}\,\tilde{S}^{\prime}(\Delta gL^{1/\nu}),$ (16)
+where $\nu$ is the correlation length exponent.
+This allows for precise determination of the critical point $g_{c}$ and exponent $\nu$ via data collapse, as demonstrated for (2+1)D O($N$) universality classes.
+2. Standard First-Order Transitions: The DRTE exhibits a characteristic discontinuity with a double-peak structure and follows the volume-law scaling
+$\frac{\partial S^{(2)}}{\partial g}=\left[2L^{d+z}\Delta f^{\prime}(g)\right]\left[w(2\beta)-w(\beta)\right]$ (17)
+where $d$ is the spatial dimension and $z$ is the dynamical exponent. The zero crossing at $g_{c}$ and the sign change of the peaks provide an unambiguous signature of phase coexistence.
+3. Weak First-Order Transitions: The DRTE obeys a hybrid scaling form
+$\frac{\partial S^{(2)}}{\partial g}=\left[2L^{d+z}\Delta f^{\prime}(g)+\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}\right]\left[w(2\beta)-w(\beta)\right],$ (18)
+which encapsulates the competition between the underlying first-order mechanism (first term) and the influence of a nearby critical fixed point (second term).
+In accessible finite sizes ($L\ll\xi$), the critical-like $L^{1/\nu}$ scaling dominates, explaining the observed data collapse with continuous exponents.
+Simultaneously, the persistent double-peak structure and zero crossing betray the first-order nature, as conclusively shown for the debated $J$-$Q$ models.
+Beyond these specific findings, the RTE/DRTE approach offers a general, unbiased, and order-parameter-agnostic diagnostic tool.
+Its applicability extends to transitions elusive to conventional local observables, including potential topological phase transitions.
+While validated here with high-precision quantum Monte Carlo, our method is readily adaptable to tensor network and other numerical techniques [87, 88].
+This work thus provides a new paradigm for characterizing quantum criticality, resolving the long-standing challenge of distinguishing weak first-order from continuous transitions in finite-size simulations.
 Acknowledgements.
-We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang and Wenan Guo.
+We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang, Zi Yang Meng, Chong Wang and Wenan Guo.
 Z.W.
 thanks the China Postdoctoral Science Foundation under Grants No.2024M752898.
-The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University.
+The work is supported by the Scientific Research Project (No.WU2025B011) and the Start-up Funding of Westlake University.
 The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd.
 for providing HPC resources.
-Supplemental Materials
 Appendix A Method
@@
 $W(J^{\prime})$ and $W(J)$ represent the weights of a same configuration but under different parameters $J^{\prime}$ and $J$.
-In this frame, we can designed the incremental process along the path composed of real physical parameters  [^1] It was virtual path before.
+In this frame, we can designed the incremental process along the path composed of real physical parameters.
 In other words, all intermediates are physical partition functions at corresponding parameters.